### v1

# Reconciling Inflation with Hubble Anisotropies

ABSTRACT
There have been persistent suggestions, based on several diverse data sets, that the cosmic expansion is not exactly isotropic. It is not easy to develop a coherent theoretical account of such a “Hubble anisotropy”, for, in standard General Relativity, intuition suggests that it contradicts the predictions of the very successful Inflationary hypothesis. We put this intuition on a firm basis, by proving that if we [a] make use of an Inflationary theory in which Inflation isotropises spatial geometry —  this, of course, includes the vast majority of such theories —  and if [b] we insist on assuming that spacetime has a strictly metric geometry (one in which the geometry is completely determined by a metric tensor), then indeed all aspects of the “Hubble field” must be isotropic. Conversely, should a Hubble anisotropy be confirmed, then either we must contrive to build anisotropy into Inflation from the outset, or we will have to accept that spacetime geometry is not strictly metric. We argue that allowing spacetime torsion to be non-zero would be by far the most natural way to accommodate such observations. Such theories make firm predictions, as for example that there should be a correlation between the degree of anisotropy at the end of Inflation and a certain specific component of the Hubble tension.
1. Inflation and the Hubble Field
Cosmic Inflation [1, 2] gives a convincing account of the formation of large-scale cosmic structure and explains the observed almost scale-invariant spectra of temperature anisotropies in the CMB; nowadays these aspects of the theory, closely linked as they are to observations [3], are regarded as the most important.
However, Inflation is still called upon to explain (or at least render less mysterious) some much more basic observations. Among these is the strange fact that, while spacetime has an observably non-trivial geometry, it seems that, at cosmological scales, three-dimensional space does not. Much effort has been expended on searches for the curvature of three-dimensional space, to no avail thus far (see [4] for a clear discussion, and [5, 6] for the latest data on this).
We are entitled to ask why our spacetime has this extraordinary structure. And it is indeed extraordinary: the only way for the spatial sections to have an effectively Euclidean [^1] We use “Euclidean” to mean “devoid of non-trivial geometry”, avoiding the word “flat” because it has misleading connotations: we mean that the curvature is zero, but also that any other measure of geometric non-triviality in three-dimensional space, such as torsion, should vanish. intrinsic geometry is if the spacetime curvature, evaluated on them, resides entirely in their extrinsic geometry. But, for technical reasons to be briefly reviewed later (see the footnote in Section 3, below), this means that our spacetime is foliated by spatial slices in a very unusual way, one not possible for a generic spacetime. This calls for an explanation. (Note that this is not the traditional “Flatness Problem”: we are referring to the highly non-generic manner in which spacetime is foliated at the present time.)
Inflation explains this strange structure by, first, supposing that the spatial curvature in the pre-Inflationary Universe was bounded; this is motivated by the standard hopes about the avoidance of singularities in some future complete theory of the earliest times. It is then plausible that the enormous, de Sitter-like expansion characteristic of Inflation reduced the region of space we can observe to such a small subspace (see [7] for this) that it becomes indeed effectively Euclidean. That sets up just the right initial conditions for the part of spacetime we can observe to evolve to its present unusual state, with spatial sections dominated by extrinsic curvature.
Notice that this argument eliminates all indications of non-trivial spatial geometry: it relies only on the assumption that the spatial sections have the structure of a differentiable manifold (and not, for example, that of a cone). That is, all classical measures of non-Euclidean spatial geometry are “inflated away” (leaving only the effects of quantum fluctuations of the inflaton, the temperature anisotropies mentioned earlier).
If indeed this picture is correct —  see below —  it is often said that Inflation predicts that all large-scale observations of the Universe (and not just the cosmic redshift field, regarded as a potentially non-constant function of direction) should, on average, reveal structure which is isotropic relative to some distinguished family of observers. One says that Inflation predicts “statistical isotropy”.
Now it should be noted that making this standard argument that “Inflation implies isotropy” reasonably rigorous is surprisingly difficult: see [8] for a good discussion, particularly of the problems associated with the quantum-mechanical aspects. And indeed, various aspects of this argument have, from time to time, been questioned. For example, it has sometimes been asked, in effect, whether Inflation really gives rise to enough expansion to “inflate away” the geometry of three-dimensional space: see for example [9] and particularly [10]. From a completely different viewpoint, it has been pointed out that the field responsible for Inflation, the “inflaton”, does not itself “inflate away”; if, therefore, the inflaton is inherently anisotropic —  if, for example, it is a vector field —  then we might not expect Inflation to predict statistical isotropy (see the discussion at the end of [8]).
However, these alternatives to the standard picture of Inflation, while interesting, seem rather contrived; in the case of “vector Inflation”, it is also difficult to arrange for the field to be stable. It seems to us that the peculiar spacetime structure we observe, with essentially all of the spacetime curvature (evaluated on spatial sections) concentrated in the extrinsic curvature of those sections, reflects the action of some profound physical process, and is no accident. We shall therefore, henceforth, cleave to the standard picture, as described above: we will take it that Inflation does indeed “inflate away” all non-trivial spatial geometry, so that the spatial sections are effectively exactly Euclidean, and therefore statistically isotropic. This approach also has the advantage of simplicity (and clarity).
All this has recently become topical and relevant, because statistical isotropy has been cast into doubt by analyses of several data sets: see for example [11, 12, 13, 14]. (See also Section 4.6 of [15] for a survey of these and other possible anomalies in the recent cosmic observations.) While this “statistical anisotropy” is by no means confirmed [16], recent observational advances promise to settle the question in the not very distant future. Since these new observations challenge the very foundations of the standard understanding of the early Universe, they certainly deserve attention.
In addition to this, the concept of “isotropy” has recently expanded and acquired a new depth; this will be an important theme in our considerations here.
Traditionally, “isotropy” meant, “isotropy of the cosmic redshift field”, since the redshifts of distant galaxies were of course the data that first revealed the expansion of the Universe. However, a novel and potentially very important aspect of the recent surge of work on statistical anisotropy is that it no longer relies exclusively on observations of redshifts. For example, [17] (see also [18, 19]) gives a discussion of the ways in which statistical anisotropy might appear in the shapes and particularly the orientations of certain very massive galactic halos. The idea here is that if somehow statistical anisotropy does occur in the initial conditions at some distinguished early time, then the anisotropy might affect the earliest phase of halo formation, and it might persist through the subsequent evolution, with potentially observable consequences. In a similar spirit, studies of galaxy clustering and alignments have also been proposed [20] as probes of large-scale anisotropy.
These remarkable new developments might well give us effective ways of studying statistical anisotropy, independently of the redshift field (which of course remains of great interest, and in which anisotropies may also exist, see for example [21] and especially the very clear presentation in [22]). (Thus the plural in the title of the present work.)
Instead of referring only to the redshift field, then, we shall speak of a Hubble field, the field representing the directed rate at which the galaxies actually separate as cosmic time passes. We will see that the two are related but far from identical in general.
All this prompts us to go back to basics and ask: is it really the case that (conventional) Inflation necessarily predicts isotropic “initial conditions” for the subsequent evolution, leading to an isotropic Hubble field?
The key point here is that, in conventional models of Inflation, the intrinsic geometry of the spatial sections certainly is “isotropised” [8]: once we accept that all traces of non-trivial spatial geometry have been eliminated, then the sections must be isotropic, simply because Euclidean space itself is indeed isotropic around any point. Thus reheating [23, 24, 25], which we can think of as the start of “well-understood physics”, does take place in an isotropic (Euclidean) three-dimensional space if conventional Inflation is valid.
Thus we are forced to re-formulate our question: does the isotropisation of the intrinsic geometry of the spatial sections at reheating imply an isotropisation of the full set of “initial conditions”, at that time?
This is not at all as obvious as it may look. The initial value problem for General Relativity requires a specification of both the intrinsic and the extrinsic geometry of the spatial sections. As we will review, the Hubble field is (primarily) a manifestation of the extrinsic geometry. The point is that it is far from clear that the extrinsic geometry of the cosmic spatial sections must necessarily behave in precisely the same way as their intrinsic geometry.
However, we have to contend with the fact that the intrinsic and extrinsic geometries are by no means entirely independent. To be more specific: the study of submanifolds in differential geometry requires us to specify two tensors on a spatial section $\Sigma$: the spatial metric $g$ and another bilinear form $h$. These two tensors are related in two ways.
First, they have to be related to each other and to the matter content of the spacetime by the initial value constraint equations [26] if they are to generate a consistent temporal evolution from $\Sigma$. This evolution endows $h$, retrospectively as it were, with a geometric meaning, as the second fundamental form of the embedding of $\Sigma$ into spacetime. Once this geometric meaning is established, it implies that there is a second, purely geometric relation between $g$ and $h$, expressed by the Weingarten equation of subspace theory (see Chapter VII of [27]).
In short, the intrinsic metric $g$ and the second fundamental form $h$ are related in various rather deep ways, both geometric and physical; so it is not out of the question that $h$ might inherit isotropy from $g$, so that the full set of “initial conditions” bequeathed by Inflation is indeed necessarily isotropic.
To summarise: the fate of the Hubble field is connected to that of the intrinsic geometry, including its isotropisation, in rather subtle ways. We wish to argue from this that it is not obvious that Inflationary isotropisation of spatial sections necessarily entails an isotropic Hubble field.
Our objective here is to understand this precisely. We will find that, in all conventional Inflationary theories (meaning those in which anisotropy is not built in from the outset), and in all theories of gravitation using strictly metric spacetime geometry (meaning geometry which is completely fixed by a semi-Riemannian metric and nothing else), the extrinsic geometry does necessarily share the symmetries of the intrinsic geometry: in all such theories, the Hubble field (and the matter fields) must isotropise if Inflation occurs. This is why it is so difficult to reconcile statistical anisotropy with Inflation.
However, we also show that this is really a consequence of the special relationship between intrinsic and extrinsic geometries in strictly metric spacetimes, and that it is not the case in the simplest and geometrically most natural extensions of GR, those in which the torsion is permitted to be non-zero [28, 29, 30].
Our principal findings regarding cosmic isotropy in torsional Inflation are as follows.
First, when torsion is present, it is indeed possible for intrinsic isotropy to co-exist with anisotropy of the Hubble field: in other words, torsion supplies a natural way of reconciling Inflation with (some) observations of statistical anisotropy.
Second, certain special kinds of torsion (the “axial” variety) make simple predictions as to the form taken by the Hubble field anisotropy: for example, in the axial case, there is always just one distinguished direction, and that direction is associated with a minimum of the magnitude of the Hubble field.
Third, the extent of the anisotropy induced by torsion is measured by parameters which appear again in the generalised Friedmann equation. Thus for example axial torsion theories imply the existence of correlation between the Hubble anisotropy and a possible torsional component of the much-discussed “Hubble tension” [31, 32, 33]; this means that observations of statistical anisotropy might in principle allow (falsifiable) predictions as to the value of that component, and vice versa.
Finally, torsion can and often does disrupt the familiar simple relation between cosmic redshift and the scale factor, leading in some cases to a very complex and subtle form of anisotropy in the redshift field.
In summary, we argue that torsion offers the most natural path to reconciling Inflation with (some) possible observations of statistical anisotropy, while linking them to other cosmological anomalies, such as the Hubble tension.
Before we begin, we should point out that we are concerned here exclusively with what might be called “geometric” anisotropies. One should bear in mind, however, the possibility of anisotropies due to other causes, such as those due to bulk flows [34, 35] or to the growth of non-linear structures [36]. These might well be just as important as geometric anisotropies, and they may be superimposed on the effects we discuss. We exclude them here only to avoid confusion. (See however Section 8, below, where we will see that this separation may not ultimately be fully tenable.)
We begin with some simple observations regarding ordinary Inflation, in the context of standard General Relativity.
2. What Inflation Does and Does Not
To reiterate: the spatial sections of our Universe, those in which (apart from small peculiar velocities) the galaxies are distributed uniformly, appear (at present, and in the directly observable past) to be Euclidean, to a good approximation [4, 5, 6]. And yet, cosmic spacetime is by no means close to being Minkowskian. Spacetime has non-trivial geometry, but space does not. Some physical process has imposed this structure on spacetime. We assert that that process is what we call “inflating away”.
In order to understand this “inflating away”, let us begin by attempting to understand the two aspects, flatness and isotropy, separately. So we begin by assuming that no anisotropy is present, and focus at first on the fate of the spatial curvature. (In this Section, we use only standard General Relativity.)
The spacetime geometry during Inflation is then approximately that of de Sitter spacetime. Since we want the spatial curvature to be non-zero initially (so that we can study what happens to it), we choose the version of de Sitter spacetime which is foliated (globally) by spacelike three-spheres.
The scale factor in this case is $L\cosh(t/L),$ where $L$ is the length scale defined by the energy density (which does not vary) and $t$ is proper time, taken to satisfy $t\,\geq\,0$ (t= 0t\,=\,0 representing the beginning of Inflation.)
The intrinsic curvature tensor is [^2] See [37] for details of the following discussion; to ease comparison with the geometry literature, we mostly follow the notations of [27].
$R(W,\,Z,\,X,\,Y)\,=\,L^{-2}\mathrm{sech}^{2}(t/L)\left(g(Y,\,Z)\,g(W,\,X)\;-\;g(X,\,Z)\,g(W,\,Y)\right),$ (1)
where $g$ is the spatial metric (that of a three-sphere here), $R$ denotes the spatial curvature tensor, and W,Z,X,YW,\,Z,\,X,\,Y are spacelike tangent vectors. This evidently takes its maximum value at the start of Inflation, and then quickly decreases towards zero: the intrinsic curvature indeed “inflates away”, exponentially quickly.
However, the Hubble parameter $H$ does not “inflate away”: just the reverse:
$H\,=\,{1\over L}\,\tanh(t/L);$ (2)
we see that the Hubble parameter increases during Inflation. This is our first indication that the Hubble field is not necessarily affected by Inflation in the same way as the intrinsic geometry of the spatial sections.
We can see the geometric significance of the $\tanh(t/L)$ factor by studying the extrinsic curvature.
The Gauss formula for submanifold curvature [27] states that the full spacetime curvature, evaluated on tangent vectors to the spatial section $\Sigma,$ is simply the sum of the intrinsic and the extrinsic curvatures, evaluated on those same vectors. The total curvature (that of de Sitter spacetime), however, is $L^{-2}\left(g^{*}(Y,\,Z)\,g^{*}(W,\,X)\;-\;g^{*}(X,\,Z)\,g^{*}(W,\,Y)\right)$, where $g^{*}$ is the spacetime metric (which can however be replaced here by the spatial metric $g$, since all of the vectors here are tangent to the spatial section); the coefficient is constant because the energy density does not change during Inflation.
Subtracting the intrinsic curvature, we see that the extrinsic curvature is given by
$E(W,\,Z,\,X,\,Y)\,=\,L^{-2}\tanh^{2}(t/L)\left(g(Y,\,Z)\,g(W,\,X)\;-\;g(X,\,Z)\,g(W,\,Y)\right).$ (3)
Far from being “inflated away”, the extrinsic curvature grows, in such a manner as to compensate for the decline of the intrinsic curvature. One might say that the extrinsic curvature has to grow in order to make this compensation possible.
The coefficient, $L^{-2}\tanh^{2}(t/L)$, is of course the square of the Hubble parameter, which is thereby revealed to be an aspect of the extrinsic geometry of the spatial sections. By the end of Inflation, the Hubble parameter has survived, indeed grown, because extrinsic geometry is immune to being “inflated away”.
Notice that the extrinsic curvature is actually zero at t= 0t\,=\,0, the beginning of Inflation [^3] One of the virtues of the “no-boundary” theory of the origin of the cosmos [38] is that it explains this initial geometry: the initial extrinsic curvature must be zero in this theory.. The situation at the start of Inflation was precisely the reverse of the one we find ourselves in today: at t= 0,t\,=\,0, the intrinsic curvature was large, while the extrinsic curvature was negligible, whereas currently the intrinsic curvature is (probably) too small to detect, while the extrinsic curvature is large, manifested in the large value of the Hubble parameter at present.
In summary: what does Inflation do? It suppresses the intrinsic geometry of the spatial sections, causing the spatial sections to have a Euclidean, isotropic structure.
What does Inflation not do? It does not suppress the extrinsic geometry of the spatial sections: in fact, just the reverse. The large Hubble field we see today is an aspect of that extrinsic geometry, and as such owes its existence to the fact that the extrinsic geometry “inflates up” instead of “inflating away”.
In view of this, we should not necessarily expect the extrinsic geometry to be isotropised by Inflation, and therefore we should not expect the Hubble field to be isotropised either. We can hope that, nevertheless, somehow it was: perhaps, in some way, the intrinsic geometry communicates its isotropisation to the Hubble field. But this is something to be proved.
In order to proceed, we need a precise definition of the Hubble field, without assuming isotropy from the outset.
3. Geometry of the Hubble Field
We now give a mathematical formulation of the Hubble field. The exposition of the geometry here is adapted to our particular needs. For alternative expositions with substantially different emphases, see [39, 13].
We begin by supposing that spacetime has a semi-Riemannian metric $g^{*}$, and a connection $\nabla^{*}$ which is compatible with $g^{*}$ but which does not necessarily have zero torsion [^4] We consistently use the asterisk superscript to denote spacetime quantities. The absence of the asterisk denotes a purely spatial quantity..
Throughout this work, we assume that spacetime can be foliated by spacelike hypersurfaces such that the galaxies are (statistically) uniformly distributed in all of them, that is, for all time. Of course, this “Copernican” condition, often called “homogeneity”, is a strong constraint; but it seems to correspond to what we observe.
Let $\Sigma$ be such a hypersurface and let $\xi$ be the corresponding local future-pointing timelike unit normal vector field. The integral curves of $\xi$ represent the distinguished observers, those who observe statistical homogeneity as above. We do not assume that they are necessarily at rest in the galaxies, nor that they are free particles.
We know that the distances between galaxies in our Universe are functions of time. Let $X$ be a vector field tangent to $\Sigma$, describing the spatial separation in $\Sigma$ of two nearby objects, which we can take to be galaxies. We call $X$ the separation field. We need a way of measuring the rate of change of these distances, as seen by the distinguished observers. We will do this using differential-geometric language, so as to obtain a formulation which can easily be adapted to the case of spacetimes with torsion.
One might think that the tangential component of $\mathcal{L}_{\xi}X$, where $\mathcal{L}$ represents the Lie derivative, would be a suitable measure of the rate of change of $X$, as seen by the distinguished observers. (Note that $\mathcal{L}_{\xi}\xi=0,$ which is reasonable since we are measuring rates of change as seen by observers with tangent vectors equal to $\xi$.) But there are two objections to it.
First, the galaxies separate for at least two distinct reasons: first, they might actually be moving systematically [^5] We will ignore the random motions of the galaxies, usually described in terms of their “peculiar velocities”. relative to the distinguished observers (this is the “bulk flow” mentioned above), and, second, the rules of three-dimensional geometry might be time-dependent. This “dynamic spatial geometry” is usually what we mean when we speak of “expanding space”, and it is this kind of time-dependent separation that concerns us here. But this second cause of changing separations is an aspect of the spacetime geometry, specifically of its metric; and that means that our measure of rates of change must likewise depend on the metric. The Lie derivative does not do so. Second, the Lie derivative is not a linear operator on vector fields over smooth functions, so it will not yield the familiar linear Hubble-Lemaître law when we apply it to the standard FRW cosmologies.
This second objection applies equally to the other obvious candidate, namely (the tangential component of) $\nabla^{*}_{\xi}X$; which does however have the virtue of depending on the spacetime geometry.
The solution to these problems is well known in differential geometry: one simply takes the difference of the two operators. This is discussed (in more formal language) in Section 4 in Chapter 6 of [40], where the deep geometric meaning of this procedure is developed. (In particular, this difference is used to map Killing vectors into the Lie algebra of the linear holonomy group of a Riemannian manifold; this is Kostant’s Theorem, Theorem 4.5 in Chapter 6 of [40].)
If we denote this operator by $\hat{\mathfrak{H}}$, then we can express the rate of separation as the vector field
$\hat{\mathfrak{H}}(X)\,=\,\tau\left(\nabla^{*}_{\xi}X\,-\,\mathcal{L}_{\xi}X\right),$ (4)
where $\tau$ denotes the tangential component.
Note that, if $F$ is a smooth function on spacetime, then the operator on the right side of this equation annihilates $F$, so indeed $\hat{\mathfrak{H}}$ acts linearly on linear combinations of scalar field multiples of spacetime vector fields; in particular, it acts linearly with respect to vector and scalar fields on $\Sigma$, when $\hat{\mathfrak{H}}$ is restricted to $\Sigma$. This is what we need: $\hat{\mathfrak{H}}$ is a linear operator which measures the rate of separation of the galaxies, taking into account the contribution of the time-dependence of the spatial geometry.
However, for this to work, $\xi$ itself must be time-independent in this sense (otherwise the “time derivative” defined by $\xi$ will be contaminated by the time variation of $\xi$), so we must impose the condition that this operator yields zero when applied to $\xi$, which means that we require [^6] This has the side-effect that the worldlines of the distinguished observers are geodesics. This may seem apt, but in fact it plays no fundamental role in our arguments from this point. Note that the worldlines of photons, to be discussed in Section 8 below, are often not geodesic in torsional theories. $\nabla_{\xi}^{*}\xi=0.$
In summary, the rate of separation is related to position by a (1, 1)(1,\,1) tensor [^7] Here, and henceforth, we indicate a (1, 1)(1,\,1) tensor by a “hat”; the metrically equivalent (0, 2)(0,\,2) tensor is denoted by the same symbol without the “hat”. on $\Sigma$. This we can call the generalised Hubble-Lemaître Law: as we will see in detail later, it reduces to the familiar Law in the cases where $\hat{\mathfrak{H}}$ (which we will call the “Hubble tensor”) reduces to a simple multiple of the identity map.
The (0, 2)(0,\,2) version of $\hat{\mathfrak{H}}$ is
$\mathfrak{H}(X,\,Y)\,=\,g(\hat{\mathfrak{H}}(X),\,Y),$ (5)
where $X$ is as above, $g$ is the spatial metric (induced on $\Sigma$ by the spacetime metric $g^{*}$), and $Y$ is any tangent vector to $\Sigma$. Note that we are not assuming that $\mathfrak{H}$ is a symmetric bilinear form, so the ordering here can be important.
Clearly $\mathfrak{H}(X,\,Y)$ represents, if $Y$ is a unit vector, the component of the rate of separation of two galaxies with separation vector $X$, in the direction of $Y$. In particular, then, $\mathfrak{H}(X,\,X)$ is a positive multiple of the radial component of the rate of separation. Having discarded the peculiar velocities, we assume that this radial component is always positive for any $X\neq 0$. The motivation here is that we want to consider a Universe which, overall, is “expanding”, in the sense that there is no direction in which it contracts or is static. This does not mean, of course, that it expands at the same rate in all directions or even that the expansion is always radial [^8] One might think that this assumption follows from the non-observation of cosmic blueshifts. However, as we will discuss in detail in Section 8 below, the redshift field does not always coincide with the Hubble field, so, strictly speaking, we cannot reason in that way..
Thus, the symmetric part of $\mathfrak{H}$ is positive-definite. Now if a real matrix has a positive-definite symmetric part, then its real eigenvalues are positive. It follows that the determinant of $\mathfrak{H}$ is positive, and hence, in particular, not zero; so $\mathfrak{H}$ is not singular.
We can understand $\mathfrak{H}$ better by splitting it using the well-known polar decomposition [41], which states that any non-singular bilinear form $B$ can be uniquely factorised as $B=Q\sqrt{B^{T}B},$ where $Q$ is an orthogonal linear map and where the square root denotes the positive square root of the positive-definite symmetric bilinear form $B^{T}B.$ This square root is itself a positive-definite symmetric bilinear form.
Applying this to $\mathfrak{H}$, we therefore write
$\mathfrak{H}\,=\,Q\,\eta,$ (6)
where $Q$ is an orthogonal transformation (in this case, since the determinant of $\mathfrak{H}$ is positive, a rotation) depending on $\mathfrak{H}$, and where [^9] The letter $\eta$ was originally the Greek equivalent of $h$, though of course it no longer is. $\eta$ is a symmetric positive-definite bilinear form or tensor on $\Sigma$, defined as the positive-definite symmetric square root of $\mathfrak{H}^{T}\,\mathfrak{H}$. Since, as we have seen, $\mathfrak{H}$ is non-singular, $\eta$ is uniquely defined.
In the case of theories based on strict metric geometry, it turns out (see below) that $\mathfrak{H}$ itself is a symmetric bilinear form. Because it is positive-definite, as discussed above, the uniqueness of the polar decomposition implies that $Q$ is the identity map in all such geometries.
In strict metric geometry, then, the generalised Hubble-Lemaître Law involves a simple symmetric bilinear form, $\mathfrak{H}$. Its geometric interpretation is given by the Spectral theorem: the eigenvectors of the corresponding linear transformation are mutually orthogonal, and the eigenvalues are positive real numbers, so the linear transformation represents a “stretching” by various amounts in three perpendicular directions. If the Hubble field is not isotropic, then we can quantify the anisotropy by studying the three real positive eigenvalues of $\mathfrak{H},$ let us call them $H_{i},\;i=1,\,2,\,3,$ which will not all be the same; they describe the rates of separation in three perpendicular directions.
Notice that, in directions other than those of the eigenvectors of $\mathfrak{H},$ the Hubble field will not be radial if it is not fully isotropic: there will be a tangential component.
If the geometry is not strictly metric, then a rotational component, measured by $Q,$ can be present; from the above discussion it is clear that it must be present if $\mathfrak{H}$ is not a symmetric bilinear form, since $\eta$ is always symmetric. In three dimensions, a rotation always defines a distinguished direction, the axis of rotation. This axis need not coincide with any of the eigenvectors of $\eta.$ In general, then, the situation is rather complicated; but note that (since every real three-dimensional linear map has at least one real eigenvector with a real eigenvalue), in all cases there is at least one direction in which the rate of separation vector is radial.
Next, we need to explain the geometric meaning of $\mathfrak{H}$.
In the strictly metric case, this is straightforward. Since, in that case, we take the torsion tensor to vanish, we have, using the definition of that tensor (and the fact that $\mathcal{L}_{\xi}X=[\xi,\,X]$, the commutator of the two vector fields regarded as derivations),
$\hat{\mathfrak{H}}(X)\,=\,\hat{h}(X),$ (7)
where $\hat{h}$ is defined by
$\hat{h}(X)\,=\,\nabla^{*}_{X}\xi.$ (8)
Evidently $\hat{h}$ tells us how the unit normal $\xi$ changes as we move around in $\Sigma$. That is, it tells us the way in which space bends into spacetime. In fact, $\hat{h}$ is the tensor on $\Sigma$ usually known [27] as the second fundamental form. Notice that $\hat{h}(X),$ for any $X$ tangential to $\Sigma,$ is itself automatically tangential to $\Sigma$: one sees this by differentiating $g^{*}(\xi,\,\xi)=-1$ and using the fact that the covariant derivative of $g^{*}$ is zero, so that $g^{*}(\xi,\,\nabla^{*}_{X}\xi)=0,$ meaning that the normal component vanishes. (That is why we did not need to take the tangential component in equation (7).)
In short, then, in the strictly metric case (only), the Hubble tensor is just the second fundamental form.
Now the second fundamental form determines the extrinsic geometry of the spacelike section $\Sigma;$ for example, it determines the extrinsic curvature tensor: we always [^10] We stress the “always”: this holds true no matter how complex the spacetime geometry may be. This is in sharp contrast with the intrinsic curvature, which has such a simple form (as in equation (1)) only in very special cases. The Gauss formula for submanifold curvature means that, if the intrinsic curvature of the spatial sections vanishes, then the full spacetime curvature, when evaluated on $\Sigma,$ is forced to take this precise, very special, form. This is what we meant by saying earlier that the flatness of the spatial sections implies that the foliation of the Universe by spacelike hypersurfaces, as we observe it now, is exceedingly non-generic. have
E(W,Z,X,Y)=h(Y,Z)h(W,X)h(X,Z)h(W,Y),E(W,\,Z,\,X,\,Y)\,=\,h(Y,\,Z)\,h(W,\,X)\;-\;h(X,\,Z)\,h(W,\,Y), (9)
for vectors W,Z,X,YW,\,Z,\,X,\,Y tangential to $\Sigma$, and where $h$ is the (0, 2)(0,\,2) version of $\hat{h}$. Note (see [37]) that this expression for the extrinsic curvature is true for any spacetime geometry endowed with a spacetime metric and a compatible linear connection: it is valid whether or not the geometry is what we are calling “strictly metric”. (That is, it is valid even if $h$ is not a symmetric bilinear form; but in that case the ordering of the vectors in equation (9) must be carefully preserved.)
As claimed earlier, then, we now see explicitly that the Hubble tensor, which in the strictly metric case being considered here coincides with the second fundamental form, is a measure of the extrinsic geometry of spacelike slices of spacetime. In principle, therefore, it is disconnected from whatever happens to the intrinsic geometry of the spatial slices; in particular, there is no reason at this point to think that the Inflationary isotropisation of the intrinsic geometry should ensure isotropy of the Hubble field, just as the “inflating away” of the intrinsic curvature does not imply any “inflating away” of the Hubble parameter —  as we saw, this does not happen. On the contrary, we expect measures of the extrinsic geometry, like $\mathfrak{H}$ and its eigenvalues, to grow during Inflation, and we should not expect these eigenvalues to grow towards a common value.
Nevertheless, particular properties of strictly metric geometry do in fact ensure that the Hubble field must share the symmetries of the spatial metric in that special case, as we now discuss.
4. Isotropisation of the Hubble Field in Strictly Metric Inflation
The description of a spatial hypersurface in spacetime involves two basic objects, the spatial metric $g$ and the second fundamental form $h$. In general, there is no particular reason to expect the two to be closely related; generically, for example, $h$ is not even a symmetric bilinear form [37] at each point.
In the specific case of strictly metric geometry —  recall that this means that the metric determines the geometry completely —  however, there is in fact a close relation: we have (see [37]) in this case, from the Weingarten equation of subspace theory [27],
$h\,=\,{1\over 2}\,\mathcal{L}_{\xi}\,g^{*},$ (10)
where $\mathcal{L}_{\xi}\,g^{*}$ is the Lie derivative (evaluated on $\Sigma$) of the full spacetime metric $g^{*}$ with respect to $\xi$, the unit timelike normal to the spatial section, and it is understood that both sides are to be evaluated on a pair of vectors tangential to that section.
Although $g^{*}$ appears in equation (10), actually the right side only expresses the “time rate of change of the spatial geometry”, that is, of the spatial metric $g$. (One can see this most easily by expressing the restriction of $g^{*}$ to $\Sigma$ in the form $\Xi\otimes\Xi\,+\,g,$ where $\Xi$ is the dual one-form to $\xi$.)
Notice that $h$ is a symmetric bilinear form in strictly metric geometry; and, since $\mathfrak{H}=h$ here, so therefore is $\mathfrak{H}.$
The spatial section $\Sigma$ is said to be isotropic if, given any pair of unit vectors at any point, there exists an isometry $\iota$ of $g$ (that is, a smooth map of $\Sigma$ to itself satisfying $\iota^{*}g=g$, where the asterisk denotes the pull-back), mapping one vector to the other.
Pick a point $p$ in a given spatial section $\Sigma$, and consider the integral curve $\gamma$ of $\xi$ passing through that point, so that $\gamma(0)=p.$ Let us suppose that $\Sigma$ is isotropic about $p$. (Of course, we do not insist on exact isotropy: all we ask is that the space near to $p$ is well-approximated by a three-dimensional space which is isotropic around $p$.)
Now, in Inflation, isotropy is approached asymptotically [8]; that is, we are not interested in a situation where isotropy holds only at a single moment in time. Instead, we want the geometry around $\gamma(t)$ also to be approximately isotropic for at least some interval, of non-zero length, of parameter values along $\gamma$. This means that the Killing vectors associated with the isotropic mappings must commute with $\mathcal{L}_{\xi}$. Therefore equation (10) implies that each isotropic mapping of $g$ is also an isotropic mapping of $h$, that is, of $\mathfrak{H},$ since the two are equal in the strictly metric case.
In particular, if we assume that Inflation has isotropised $g$, then it also must isotropise $\mathfrak{H}$. In strictly metric geometry, Inflationary isotropisation of the intrinsic geometry necessarily entails isotropisation of the Hubble field.
Since $\mathfrak{H}$ is a symmetric bilinear form here, and since the isotropy group acts irreducibly, Schur’s lemma (see Appendix 5 of [40]) implies that, if the Hubble field is isotropic, then $\mathfrak{H}$ must be a multiple (which, since we are assuming that the galaxies are always uniformly distributed, must not depend on position in $\Sigma$) of the spatial metric tensor $g$. That multiple is of course the Hubble parameter $H$, the thrice repeated positive eigenvalue of $\mathfrak{H}:$
$\mathfrak{H}\,=\,H\,g.$ (11)
Since both $\mathfrak{H}$ and $g$ are isotropic, so is $H$, in the obvious sense. Thus $H$ can at most be a function of time.
Of course, equation (11) means that $\hat{\mathfrak{H}}$ acts simply by multiplication by $H$. Thus we see that our proposed definition of the Hubble field gives us a generalised Hubble-Lemaître law which does indeed reduce to the classical law when we return to spatially isotropic General Relativistic cosmologies.
In the simplest cases one can easily confirm that $H$ is as expected. If we choose a (proper) time coordinate such that the unit normal is $\xi=\partial_{t},$ and set $g=a(t)^{2}\,f,$ where $a(t)$ is the scale factor and $f$ is a time-independent isotropic three-dimensional spatial metric tensor, then we have
$\mathfrak{H}\,=\,{\dot{a}\over a}\,g,$ (12)
where the dot denotes differentiation with respect to proper time. Thus we regain the familiar expression $\dot{a}/a$ for the Hubble parameter in this case, and we see explicitly that $\mathfrak{H}$ inherits isotropy from $g$.
In the case of the de Sitter spacetime discussed above, we have
$\mathfrak{H}\,=\,L^{-1}\tanh(t/L)\,g,$ (13)
where $g$ is the standard metric on the three-sphere, and again we see that $\mathfrak{H}$ is isotropic because $g$ is so. Notice also that, as expected, during Inflation all measures of the extrinsic geometry, including the Hubble tensor/parameter, increase as the intrinsic geometry approaches Euclidean geometry.
Inflation of a strictly metric spacetime, then, is a process whereby the Hubble tensor grows (from a very small or zero initial value) to a tensor which, by the time reheating begins, has a large magnitude and is as isotropic as the spatial metric.
There are two remarkable aspects of this argument. The first is that we have used the word “Inflation” in its broadest possible sense: as referring to an early phase of explosive expansion. The argument therefore applies to all of the many conventional (scalar) Inflationary models, and so the conclusion cannot be avoided by tinkering with the details (the precise nature of the inflaton potential or other features of its Lagrangian, and so on).
Second, and still more striking: the remainder of the argument is purely geometric —  it makes no reference whatever to the field equations for gravity. It depends only on the assumption that the geometry of spacetime is strictly metric, an assumption that, without further input, leads inevitably to equation (10). Thus, modifying the matter content or changing the Einstein-Hilbert action will make not the slightest difference to the argument, as long as we use strictly metric geometry.
We thus reach a very strong conclusion: a confirmed observation of statistical anisotropy would mean either that
[a] All conventional models of Inflation —  that is, all Inflationary models which induce isotropisation of the intrinsic spatial geometry —  are seriously defective,
or that
[b] The geometry of spacetime is not strictly metric.
We wish to argue that option [a] should be rejected: the successes of conventional Inflation are too numerous and various. Therefore option [b] would be necessary if statistical anisotropy were to be confirmed observationally. The question is whether it is sufficient: does a non-strictly metric spacetime geometry permit a statistical anisotropy of some kind, even if Inflation has isotropised the intrinsic spatial geometry?
We will now argue that it does, and that in fact it does so in an extremely natural manner. Whether it can do so in a manner consistent with (all of) the observations is of course another matter. In fact, we will see that option [b] does permit some kinds of anisotropy, but, at least in the concrete examples we will study, not all.
5. Isotropisation and Torsion: General Theory
The familiar assumption, in standard General Relativity, that the spacetime torsion should vanish as a matter of principle, is almost impossible to justify [^11] On the other hand, we feel that the assumption that the linear connection is a metric connection [42] is physically justified in most cases (the exception being the special case in which the connection is non-metric yet “integrable”, so that lengths of four-vectors are still globally well-defined, as they appear to be in our Universe —  particle masses are just lengths of momentum four-vectors after all). We always assume that our connections are metric connections. The reader who disagrees with this restriction can interpret it as a mere simplifying assumption; and it is certainly true that allowing non-metricity would complicate our discussion very substantially. [28, 37]. This is particularly clear in the case of cosmological spacetimes.
The idea that the linear connections of cosmological spacetimes might have non-zero torsion has a surprisingly long history: see [43, 44]. (See [45, 46] for general discussions of torsional cosmology.) Current interest in spacetime torsion focuses on recent observational developments; for example, it has been suggested that torsion might be relevant to several of the cosmological “tensions” [47, 48, 49], particularly the notorious “Hubble tension”. Recently, the very interesting possibility that torsion might play an important role during Inflation has also been explored; see for example [50, 51].
When one proposes to investigate torsional theories, one has to face the fact that there is a very large number of such theories, and unfortunately it has to be said that, at present, none of them is strongly favoured, observationally, over the others. We will deal with this by remaining at the maximal level of generality. (Later we will focus on the simplest such theory, the Einstein-Cartan theory, and on the simplest variety of torsion, axial torsion, but we do this so as to make a point in the clearest possible manner, not because we are committed to these assumptions as a matter of principle.)
We characterised conventional Inflation as a process which explains, among many other things, the fact that we do not observe any evidence for non-Euclidean spatial geometry. That included intrinsic spatial torsion (and its derivatives [37]).
Thus we take it henceforth that, by the end of Inflation, the spatial sections have no non-trivial geometry of any kind: they are therefore completely isotropic. That does not imply that the full spacetime torsion must vanish, any more than the flatness of the spatial sections implies that cosmic spacetime is Minkowskian in conventional Inflation.
In the preceding Section we saw that, in the absence of torsion, spatial isotropisation necessarily entails an isotropic Hubble field. Now let us see what happens when torsion is allowed to be non-zero.
We immediately see that non-zero torsion affects our entire discussion in the previous section, because already we used the zero-torsion condition in moving from equation (4) to equation (7). Using the definition of spacetime torsion, $T^{*}(A,\,B)=\nabla^{*}_{A}B-\nabla^{*}_{B}A-[A,\,B]$ for arbitrary spacetime tangent vectors A,B,A,\,B, we now have
$\hat{\mathfrak{H}}(X)\,=\,\hat{h}(X)\,+\,\tau\,T^{*}(\xi,\,X),$ (14)
where $X$ is tangential to $\Sigma$, where $\hat{h}$ is still the second fundamental form (as given in equation (8)), and where, as before, $\tau$ means that the component tangential to $\Sigma$ has been taken. (The same proof as before shows that $\hat{h}(X)$ is automatically tangential to $\Sigma$, but that is not necessarily true of $T^{*}(\xi,\,X),$ so we have to take the tangential component explicitly.)
[…]
$T^{B}(X)\,=\,T^{*}(\xi,\,X),$ (15)
where $X$ is tangential to $\Sigma$ (but note that $T^{B}(X)$ can have both tangential and normal components).
The geometric meaning of the bulk torsion is not very clear at this point. It will turn out that it captures those parts of the spacetime torsion which are neither intrinsic nor extrinsic, that is, those parts which do not control the torsional geometry either within a spatial section $\Sigma$ (the intrinsic torsion), nor in its immediate neighbourhood (the extrinsic torsion). Instead $T^{B}$ is relevant to the effects of torsion on photons arriving here and now from the deeps of cosmological spacetime —  from the “bulk” of spacetime. Thus it is important here only as it affects the redshift field, so we postpone discussing it until Section 8 below, which is concerned with that aspect of observable anisotropy.
It will be convenient to decompose the bulk torsion into a part that is tangential to $\Sigma$, together with a part normal to it. We define a (0, 2)(0,\,2) tensor $\beta$ on $\Sigma$ by
$\beta(X,\,Y)\,=\,2\,g\left(\tau T^{*}(\xi,\,X),\,Y\right),$ (16)
the factor of 2 being included for later convenience. In terms of the (0, 2)(0,\,2) versions, we have now from equation (14),
$\mathfrak{H}\,=\,h\,+\,{1\over 2}\beta.$ (17)
Note that $\beta$ has no particular symmetry property, as a bilinear form, in general.
The normal component of $T^{B}$ is measured by a one-form $b$ on $\Sigma$ defined by
$b(X)\,=\,2g^{*}(\xi,\,T^{*}(\xi,\,X));$ (18)
it does not affect the Hubble tensor, but it will reappear in Section 8.
Clearly $\mathfrak{H}$ need not be a symmetric bilinear form here, but the relevant parts of our previous discussion continue to apply, since we did not assume symmetry there. Motivated by the absence of cosmic blueshifts, we continue to assume that the symmetric part of $\mathfrak{H}$ is positive-definite, whence we conclude that the matrices of $\mathfrak{H}$ are non-singular and therefore have a unique polar decomposition, allowing us to define a unique $\eta$ tensor, constructed from $\hat{\mathfrak{H}}$, just as before:
$\eta\,=\,\sqrt{\mathfrak{H}^{T}\mathfrak{H}}.$ (19)
The torsional generalisation of equation (10) was found, by suitably generalising the Weingarten equation, in [37]:
$h(X,\,Y)\,=\,{1\over 2}\,\mathcal{L}_{\xi}\,g^{*}(X,\,Y)\,+\,g\left(\tau K^{*}(X,\,\xi),\,Y\right),$ (20)
where as usual X,YX,\,Y are tangential to $\Sigma$, $\mathcal{L}_{\xi}\,g^{*}$ is as before (again to be understood as being restricted to $\Sigma$), and where $K^{*}$ is the spacetime contortion tensor, which is defined as the difference between $\nabla^{*}$ and $\nabla^{0},$ the fictitious connection that spacetime would have had if the torsion had been zero. (In other words, the contortion measures the extent to which the geometry deviates from being strictly metric.) Notice that equation (20) implies the fact, mentioned earlier, that $h$ is symmetric when torsion is absent.
One can show that $g^{*}\left(K^{*}(A,\,\xi),\,B\right)$ is always antisymmetric in $A$ and $B$; from this it follows that the second term in equation (20), $g\left(\tau K^{*}(X,\,\xi),\,Y\right)$, defines a two-form on $\Sigma:$
$\varkappa(X,\,Y)\,=\,2\,g\left(\tau K^{*}(X,\,\xi),\,Y\right).$ (21)
We therefore have a decomposition of the second fundamental form into symmetric and antisymmetric parts: from equations (20) and (21),
$h\,=\,{1\over 2}\,\mathcal{L}_{\xi}\,g^{*}\,+\,{1\over 2}\,\varkappa.$ (22)
Clearly $\varkappa$ is the part of the torsion that contributes to the extrinsic geometry of $\Sigma,$ as measured by $h$. We therefore call it the extrinsic torsion form. Evidently $h$ is not a symmetric bilinear form if the extrinsic torsion form is not zero.
Equation (22) is the fundamental geometric relation here. It tells us that the embedding of a spacelike hypersurface into spacetime is determined by the time rate of change of the intrinsic spatial metric, together an additional “twisting of space into spacetime” measured by the extrinsic torsion form; this is the geometric meaning of $\varkappa$.
Combining equations (14) and (20), we find
$\mathfrak{H}(X,\,Y)\,=\,{1\over 2}\,\mathcal{L}_{\xi}\,g^{*}(X,\,Y)\,+\,g\left(\tau K^{*}(X,\,\xi),\,Y\right)\,+\,g(\tau\,T^{*}(\xi,\,X),\,Y),$ (23)
which, using the definitions of $\varkappa$ and $\beta$, can be written simply as
$\mathfrak{H}\,=\,{1\over 2}\,\mathcal{L}_{\xi}\,g^{*}\,+\,{1\over 2}\,\varkappa\,+\,{1\over 2}\,\beta,$ (24)
where, as ever, it is understood that all of these objects are to be evaluated on pairs of vectors tangential to $\Sigma$. The $\eta$ tensor is constructed from this, as usual, according to equation (19).
Like $h$, $\mathfrak{H}$ is not, generically [^12] One can contrive a situation in which the antisymmetric part of $\beta$ exactly cancels $\varkappa$, so that $\mathfrak{H}$ becomes symmetric. Clearly this is a very special, artificial arrangement, and the reader can readily verify that it cannot happen in the cases of physical interest we will study below., a symmetric bilinear form when the torsion is not zero. We saw earlier (Section 3, above) that this means that a non-trivial rotational component, $Q$ in equation (6), must be present. In short, a non-trivial rotational component of $\mathfrak{H}$ is a signal that torsion is present. This is in fact the simplest way of understanding the geometric significance of torsion in cosmology.
It follows that not every form of anisotropy can be realised by torsion. For example, the kind of theoretical anisotropy we discussed in the purely metric case (which is possible in the absence of Inflation), where there is radial separation (at three different rates) along three mutually perpendicular axes (the eigenvectors of $\mathfrak{H}$ when it is a symmetric bilinear form) is impossible in torsional spacetime geometry. We will discuss this in more detail later, in Section 6.
Now let us focus on the Inflationary era, and let us follow the discussion of the previous section as closely as possible. The guiding principle here is that, after all, the absence of torsion from that discussion was incidental (or, better, an historical accident [28]).
We agreed earlier to accept that Inflation erases all forms of non-trivial intrinsic geometry in $\Sigma$, including the intrinsic torsion (the torsion of the induced linear connection on $\Sigma$). By the same logic, we take it that the measures of the extrinsic geometry of the spatial sections have increased.
More specifically, we mentioned earlier that, in the “no-boundary” theory [38] of the earliest Universe, $h=0$ at the “beginning of time”; and this is indeed a natural condition to impose at the start of Inflation, whether or not one accepts the “no-boundary” theory. For otherwise we would have to explain why $h$ takes some specific form at the beginning, and this mean importing a large number of seemingly arbitrary parameters into the theory.
The same reasoning applies here, and so, if we accept this line of argument, then both terms on the right of equation (22) must (independently) vanish at that time. Both terms will however grow as Inflation proceeds, so, in particular, the extrinsic torsion $\varkappa$ will not be negligible by the time of reheating: as we will see in detail later, it must grow, to compensate for the loss of the intrinsic torsion.
As before, the spatial metric $g$ will, with these assumptions, be isotropised by Inflation. Then the first term on the right of equations (22) and (24) isotropises, by the same argument as in the strictly metric case. But does this mean that $h$ and $\mathfrak{H}$ also isotropise?
We have no reason to think so; $\varkappa$ and $\beta$ are not controlled by the intrinsic geometry. In fact it can be shown (Appendix 5 of [40]) that a real antisymmetric bilinear form like $\varkappa$ can be isotropic, in odd numbers of dimensions (3 in our case here), only if it is zero. Thus we see that $h$ can be isotropic only if the extrinsic torsion vanishes. In other words, the extrinsic geometry is automatically anisotropic if the extrinsic torsion does not vanish.
Similarly, we have no reason to think that $\beta$ will isotropise; it certainly does not do so in the concrete examples we will consider [^13] One can however ask what happens if we force a torsional geometry to be highly symmetric. This question has led to important classification theorems: see [52, 53, 54, 55]. Our point here, however, is just that Inflation alone does not impose such a high degree of symmetry on spacetime..
In short: the Hubble field is always isotropic in a zero-torsion Inflationary theory which isotropises the spatial metric; but (unless we force it to be so, “by hand”) it is never isotropic in an inflationary theory with non-vanishing torsion.
Equation (24) illustrates the situation very clearly. In a purely metric spacetime, the isotropisation of the spatial metric isotropises the Hubble field. In the presence of torsion, however, there are two distinct possible sources of anisotropy. First, the spatial sections will, if $\varkappa\neq 0,$ embed into spacetime in an anisotropic way, specified by the detailed structure of $\varkappa$. Secondly, if $\beta\neq 0,$ anisotropies in the “bulk” (regions of spacetime remote from us) can give rise to anisotropic effects in signals we observe, depending on the properties of $\beta$.
This is why it is natural to propose that torsion may help us to reconcile Inflation with purported statistical anisotropies of various kinds: torsion provides us with a large (though, as we saw, not limitless) repertoire of techniques to induce anisotropy in those parts of the Hubble field not controlled by the intrinsic geometry of spatial sections.
Now, by the familiar argument using Schur’s lemma, when the spatial metric $g$ is isotropic, the first term on the right in equation (22) is a position-independent, isotropic multiple of $g$, which we continue to call $H$, the Hubble parameter:
$h\,=\,Hg\,+\,{1\over 2}\,\varkappa.$ (25)
We see now that the Hubble parameter mixes with the extrinsic torsion. The two terms on the right in equation (25) are of equal standing, and we should expect both to be crucial determinants of the geometry and physics at the end of Inflation.
It is interesting that, since $\varkappa$ is antisymmetric, the Hubble parameter is one third of the trace of the second fundamental form, which is usually called the mean curvature [^14] By an odd coincidence, the mean curvature is usually denoted $H$ in the differential geometry literature.; this is true whether or not torsion is present. That explains the geometric meaning of $H$ when the torsion is not zero.
Applying Schur’s lemma to equation (24) we have, in the spatially isotropic case,
$\mathfrak{H}\,=\,Hg\,+\,{1\over 2}\,\varkappa\,+\,{1\over 2}\,\beta.$ (26)
The consequences of having an anisotropic Hubble tensor is most clearly explained if we focus on torsion (and contortion) tensors of a special kind. We stress that this is not really necessary: we focus on these special cases mainly because, in them, the analysis of the Hubble tensor can be made explicit and complete. Let us see how this works.
6. Example: Dirac Field as Source of Torsion
Clearly, if we want to use torsion to reconcile a possible statistical anisotropy with Inflation, we need a source for torsion. As is well known [28], spacetime torsion is associated, in the many theories that incorporate it, with the presence of some kind of fermionic matter, that is, matter governed by the Dirac equation. (For theories of “spinor Inflation”, see [56] and papers citing it.)
In the study of the self-adjoint Dirac operator acting on fermionic fields, it is natural [57, 58] to focus attention on “axial” torsion tensors [^15] The idea that axial torsion might be connected with Inflation has been discussed elsewhere, see for example [59]., those which can be represented as three-forms; that is, those which (in their (0, 3)(0,\,3) versions) are antisymmetric in every pair of “slots”. This is the natural way to adapt the geometry to the (Clifford) algebra of the Dirac $\gamma$ matrices. This argument applies in principle to any torsional theory, though of course one has to examine the consequences, via the field equations of any specific theory, for the spin density tensor; see the next Section for the case of the Einstein-Cartan theory.
In this Section we discuss anisotropies due to torsion tensors of this kind, in detail.
Using the definition of the contortion tensor, one finds that, in the axial case, the one-form $b$ defined in equation (18) is zero, while $\beta$ (see equation (16)) is forced to be an antisymmetric bilinear form, and in fact
$\beta\;=\;-\,2\varkappa\,.$ (27)
When the torsion is axial, then, the bulk torsion is completely controlled by the extrinsic torsion; for example, the bulk torsion has to be anisotropic in this case. (Recall that the extrinsic torsion must be anisotropic under all circumstances if it does not vanish, so the existence of some kind of anisotropy is a prediction of the theory.)
When the three-dimensional metric is isotropic, we now have, from equation (26),
$\mathfrak{H}\,=\,Hg\,-\,{1\over 2}\varkappa\,$, (28)
where $H$ is as usual, for axial torsion. Compare this with equation (25) (which is valid whether the torsion is axial or not): $h$ and $\mathfrak{H}$ are similar but not the same when the torsion is axial.
Now, since $\varkappa$ is antisymmetric, we have in the axial case
$\mathfrak{H}^{T}\mathfrak{H}\,=\,H^{2}g\,-\,{1\over 4}\varkappa^{2}.$ (29)
The matrix of $\varkappa$ is a real antisymmetric $3\times 3$ matrix, so, if it is not zero, it has one eigenvector with zero eigenvalue, and two other eigenvalues which are pure imaginary and mutually conjugate. It follows that there exists a real orthonormal basis, with the zero-eigenvalue eigenvector as the third member, with respect to which the matrix of $\varkappa$ is
$\varkappa]\,=\,\left(\begin{array}[]{ccc}0&\kappa&0\\ -\kappa&0&0\\ 0&0&0\\ \end{array}\right),$ (30)
where $\kappa$ is a real quantity, which may not be a constant in either direction or time, as we will discuss. In this basis, $\varkappa^{2}$ has a diagonal matrix with entries $\kappa^{2},\,-\kappa^{2},\,0,$ and so the matrix of the $\eta$ tensor (the square root of $\mathfrak{H}^{T}\mathfrak{H}$) is, from equation (29),
$\eta]\,=\,\left(\begin{array}[]{ccc}\sqrt{H^{2}+\kappa^{2}/4}&0&0\\ 0&\sqrt{H^{2}+\kappa^{2}/4}&0\\ 0&0&H\\ \end{array}\right).$ (31)
This is the predicted form of the $\eta$ tensor as Inflation ends and reheating [23, 24, 25] begins.
We see that $\varkappa$ defines a distinguished direction, the direction of its eigenvector with zero eigenvalue; this is also the direction in which $\eta$ has an eigenvector with eigenvalue $H$. Furthermore, it is not difficult to show that the rotation transformation, $Q,$ in this case has a matrix, relative to this basis, of the form
$Q]\,=\,\left(\begin{array}[]{ccc}{H\over\sqrt{H^{2}+\kappa^{2}/4}}&{-\kappa/2\over\sqrt{H^{2}+\kappa^{2}/4}}&0\\ {\kappa/2\over\sqrt{H^{2}+\kappa^{2}/4}}&{H\over\sqrt{H^{2}+\kappa^{2}/4}}&0\\ 0&0&1\\ \end{array}\right).$ (32)
As we foresaw, this rotation is always non-trivial if the extrinsic torsion does not vanish. This “Hubble twist” is characteristic of the presence of extrinsic torsion.
The rotation is clearly around an axis which coincides with the distinguished axis of the $\eta$ tensor. (As mentioned earlier, this does not happen in general, so this is an indication that this very simple special case may in fact be relevant to the observations, which do indeed suggest the existence of a single distinguished axis.) This distinguished direction is the unique real eigenvector of the Hubble tensor, the eigenvalue being of course $H$. We stress that the rate of separation vector is never radial except in the distinguished direction, even though the spatial geometry is isotropic.
Given that the anisotropy itself is barely observable if at all, the reader may well question whether transverse “expansion” of this kind is observable. However, in the case of peculiar velocities, it is well known that this can in fact be done in principle, though in practice it is difficult: one makes use of the transverse Doppler effect [60, 61]. Of course, the cosmological redshift is not a Doppler effect, but it is reasonable to suppose that it might be possible to adapt the methods of [60, 61] to the cosmological case. This remains to be seen. (However, the interpretation of variations in the redshift field can be a formidably complex matter, even when only standard physics is in play: see [62].)
Clearly $\kappa$ controls all aspects of the anisotropy in the Hubble field, just as $H$ controls the overall expansion. Let us define a “Hubble eccentricity” $\varepsilon$ by
$\varepsilon\,=\,{\kappa\over H};$ (33)
This gives us a dimensionless measure of the anisotropy; it is also a measure of the relative importance of extrinsic torsion to extrinsic curvature.
In terms of $\varepsilon,$ the angle of the “Hubble twist” is given by
$\psi\,=\,\arccos\left({1\over\sqrt{1+\varepsilon^{2}/4}}\right).$ (34)
We saw that the algebra here distinguishes one particular direction. It is natural (though this cannot be deduced formally from anything we have said) to assume that $\kappa$ itself is also symmetric about this axis. That is, if it is not a constant, it is a function only of the spherical polar angle $\theta$ away from this axis (and of time), and does not depend on the azimuthal angle. With this assumption, the entire system is completely symmetric about this direction. In other words, the geometry is not fully isotropic, but it is still partially isotropic in the planes perpendicular to the distinguished direction.
We stress again that we have focussed on axial torsion because of its simplicity, not because we are convinced that this is necessarily the form that torsion takes during Inflation. But it is instructive, in the sense that it clarifies the kinds of effects to which torsion can give rise.
We can see four facts very plainly, in the case of axial torsional theories of cosmic anisotropy.
[a] The predicted anisotropy singles out a distinguished direction; the Hubble parameter $H$ is the rate of separation in that direction. The rates of separation in other directions are still controlled by $H,$ so we can say that the Hubble parameter is still a measure of the overall rate of separation of the galaxies. It coincides with the “Hubble parameter” in discussions where the anisotropy is neglected, as for example in the study of the “Hubble tension”.
[b] We saw earlier that not every form of anisotropy can be explained by extrinsic torsion, because torsional anisotropy always has a rotational component. In the axial case, we can see this very explicitly: the parameter $\varepsilon$ that governs the magnitude of the anisotropy is the same parameter that determines the Hubble twist. Given one, we can predict the other unequivocally, and it is not possible to have one without the other. The rotational component may be directly observable, through a cosmological analogue of the transverse Doppler effect, though whether this is really possible remains to be confirmed.
[c] The magnitude of the Hubble field always has its minimum in that distinguished direction. The location of the maxima depends on the form of $\kappa$ as a function of $\theta$; in the special case where $\kappa$ is independent of $\theta$, they occur in all directions at right angles to the distinguished direction, and the ratio of the maximum magnitude of the rate of separation vector to its minimum is $\sqrt{1+{\varepsilon^{2}\over 4}}.$
[d] The Hubble field is isotropic in the planes perpendicular to the distinguished direction: that is, the isotropy group has been “broken” by torsion, but part of it still survives. The rotation $Q$ is in this case just a rotation in those isotropic planes.
This, then, is the predicted form of the geometric anisotropy at the end of Inflation. This geometric anisotropy would have left traces on the matter distribution at reheating. We now ask exactly how that works; the hope being that, by working backwards from current observations, one will obtain a theory that can be falsified by deductions potentially in conflict with [a], [b], [c], or [d].
7. How Torsional Inflation Sets up Initial Conditions for Reheating
The Inflationary era establishes a certain set of “initial” conditions at the time when reheating starts: “initial” in the sense that reheating is the era when the first particles [23, 24, 25] came into existence, so that thenceforth we (probably) have well-understood physics that can be used to predict the further evolution, up to the present, with some confidence. By contrast, the detailed physics of the Inflationary era itself is much less clear; so the end of that era is an interesting time for us to consider.
We assume that the inflaton carries both an energy density and a spin density. The inflaton itself does not “inflate away,” so nor do these densities. We saw earlier (Section 2) the consequence of this in the case of energy density: because the inflaton energy density does not decrease, the decrease in the intrinsic curvature had to be compensated by an increase in the extrinsic curvature. We need to argue that the analogous process occurs for the extrinsic torsion: that is, when the intrinsic torsion has “inflated away”, we should find that the extrinsic torsion has increased so that, by the end of Inflation, it alone completely accounts for the spin density, which remains (approximately) unchanged. Our first objective is to show that this is the case.
We also need to understand in detail how torsion imposes anisotropy on the “initial” physics. The anisotropy then propagates to potentially observable traces at later times: see the discussion of this in [17].
In General Relativity, initial conditions involve the spatial metric $g$ on a hypersurface $\Sigma,$ along with a (0, 2)(0,\,2) tensor $h$; the two are related to each other and to the energy density [^16] We always absorb the cosmological constant into this quantity. $\rho$ by one of the familiar constraint equations:
$16\,\pi\,G\,\rho\,=\,S\,+\,\left(\mathrm{Tr}\,\hat{h}\right)^{2}\,-\,\mathrm{Tr}\,\hat{h}^{2},$ (35)
where $G$ is as usual, $S$ is the three-dimensional scalar curvature associated with $g$, $\hat{h}$ is as in Section 3 above, and $\mathrm{Tr}$ denotes the trace map. This equation is nothing but the familiar General Relativistic Friedmann equation, as can readily be seen by substituting from equation (25) with $\varkappa=0.$
We will show that the role of this equation is to communicate the symmetries of the geometry to symmetries of the physics.
The scalar curvature $S$ automatically inherits any symmetry of the spatial metric; it follows that, in General Relativity, the symmetries of the energy density, if any, are dictated by the symmetries of the extrinsic geometry, since the remaining terms on the right side of equation (35) are determined by $\hat{h}$. However, we know that, in this case, the extrinsic geometry inherits isotropy from the intrinsic geometry. So any symmetry of $g$ is inherited by $\rho.$ In particular, this is why the “initial conditions” at the start of reheating include a matter distribution which is necessarily isotropic, in the case of conventional Inflation in General Relativity.
Now let us generalise this whole discussion to a torsional theory. The simplest of these by far is the Einstein-Cartan theory [28], which, as has been argued very plausibly, is what we would call “General Relativity” if historical circumstances had been slightly different. It is distinguished by having a non-propagating torsion tensor, which may be too simple for the most general applications; but, for applications to Inflation, in which the matter content continuously permeates the entire Universe at reheating, this is less of a concern, so we adopt this model for convenience. In any case, our discussion can be adapted to torsional theories with more complex structures.
The field equations are
$Ric^{*}\,-\,{1\over 2}\,S^{*}\,g^{*}\,=\,8\pi G\,P^{*},$ (36)
and, for axial torsion,
[…]
Since $\sigma^{*}$ is a three-form, it is useful [57] to replace it with its spacetime Hodge dual, which is a one-form $\varsigma^{*}$ containing the same information. In particular, the spatial components of $\varsigma^{*}$ are the spin densities.
[…]
From the form of the matrix given in equation (30), and using the field equation (37), we see that two of the spin density components, $\varsigma^{*}_{1}$ and $\varsigma^{*}_{2}$, are forced to be zero. But the third spin density, $\varsigma^{*}_{3}$, is not. This spin density, the one along the distinguished direction, has not inflated away: it is the spin density of the inflaton, which is immune to Inflation. Using equation (30), we see in fact that this surviving spin density is proportional to the extrinsic torsion parameter, $\kappa:$
$\kappa\,=\,8\pi G\,\varsigma_{3}^{*}.$ (38)
As we hoped, Inflation has preserved the spin density of the inflaton, by transferring the intrinsic torsion to the extrinsic torsion, just as it preserves the energy density by transferring the intrinsic curvature to the extrinsic curvature, measured by the Hubble parameter. In axial torsion theories, then, $\kappa$ has a physical interpretation as the usual multiple of the component of the spin density in the distinguished direction.
Note carefully that nothing we have said requires $\kappa$ to be a fully isotropic function: it need only be symmetric in the plane perpendicular to the distinguished direction. Equation (38) forces the spin density to have the same symmetry, but no more. That is, we should expect that both $\kappa$ and the spin density will be constant multiples of a non-trivial periodic function of $\theta,$ the angle between a given direction and the distinguished direction. The inflaton field is not fully isotropic at the end of Inflation, and nor is the spin density.
Turning now to the energy density: we need the torsional analogue of the Friedmann equation. As in the case of General Relativity, this equation is really nothing more than one of the constraints which have to be satisfied by the “initial” data. It was given in [37], to which we refer the reader:
[…]
here, as before, $\rho$ is the energy density, but now $S$ is the scalar curvature of the full linear connection, including spatial torsion and its derivatives, on the spatial section. Note that, in the presence of (extrinsic) torsion, $h$ is not a symmetric bilinear form, so the final term is not the square of $\hat{h}$; this proves to be important (see [37]).
Since we are assuming that all geometric quantities on the spatial section have inflated away by the end of Inflation, we set $S=0$. Having done this, we see once again that the (generalised) Friedmann equation only involves $\hat{h}$, that is, only the extrinsic geometry. It will however involve all contributions to $\hat{h}$, including the extrinsic torsion. Thus we see that the extrinsic torsion determines the extent to which anisotropies are built into the “initial conditions”.
Returning to equation (39): we see that the torsion not only controls the “initial” anisotropies: it also shapes the form and the magnitude of $\rho$. Let us see this in detail.
Substituting from equation (25), and using the expression for $\varkappa$ given in equation (30), one finds that the generalised Friedmann equation, evaluated at the end of Inflation, takes the form
$8\pi G\,\rho\,=\,3\,H^{2}\,-\,{1\over 4}\,\kappa^{2}.$ (40)
Notice that this equation has not been derived by assuming a FRW form for the spacetime metric: it results simply from a combination of spatial isotropy and the appropriate constraint equation.
There are two points to notice here. The Hubble parameter itself is always isotropic, even in the presence of torsion. But we saw earlier that $\kappa$ is not necessarily an isotropic function: we decided that it is natural to assume that it is cylindrically symmetric around a distinguished direction. Equation (40) therefore tells us that the energy density at the time when reheating is about to start is also not isotropic. The anisotropy of the extrinsic geometry leaves its mark on the “initial” energy density, even though the underlying space is isotropic; and this mark may persist to recent times, as the system evolves.
One might try, for example, to use this general idea to develop a theory of the initial conditions leading to the anisotropies which may have been observed in massive galactic haloes [17, 18, 19, 20], independently of whatever one observes in the redshift field (which we about to discuss in the next Section). We do not claim that a realistic account of these haloes can be constructed using this simple example. The point is that the necessary, no doubt complicated, initial conditions for this kind of anisotropy can be can perhaps be constructed by focusing on the extrinsic torsion, and that this cannot be done in a strictly metric context for Inflation. (In this connection, see [63] for the important proof that the torsional constraint equations are preserved by the time evolution.)
The second point is that, if we use $\rho$ and equation (40) to reconstruct $H$ at very early times, we will under-estimate it if we are not aware that torsion is present. The extent of this under-estimation is of course measured by $\kappa$, which however is also a measure of the degree of anisotropy in the Hubble field. It is useful to use equation (33) to write equation (40) in terms of the Hubble parameter and the “Hubble eccentricity” $\varepsilon$:
$8\pi G\,\rho\,=\,3\,H^{2}\,\left(1\,-\,{\varepsilon^{2}\over 12}\right).$ (41)
In principle, therefore, a measurement of the Hubble anisotropy will directly correlate with the apparent reduction in the value of $H$, simply because the same parameter controls both effects. (We might also find that the extent of the reduction will depend on direction, again in a way that would correlate with observations of the detailed structure of the anisotropy.)
As the Universe evolves towards the present, $\rho$ and $H$ decrease, and the approximate correctness of the FRW model with Euclidean spatial sections tells us that $8\pi G\,\rho$ approximates $3\,H^{2},$ so equation (40) tells us that $\kappa$ must decrease and become so small that it is difficult to detect at the present time. The upshot will be that both the anisotropy in $\rho$ and the apparent reduction of $H$ by $\kappa$ will be hard to detect in recent times, but could well be significantly more marked at early times. Thus in particular we should expect the value of $H$ deduced from data on the early Universe to appear to be smaller than in observations of the relatively more recent cosmos. If we are very fortunate we might also find that this early/late disparity depends on direction.
Of course, we do not claim that such a simplistic analysis can solve the famed “Hubble tension”. Probably the tension is the result of several contributing factors [32]. But it may be that torsion, quantified by $\kappa$, is one of those factors. When the other contributions to the “tension” have been fully accounted for, we may find that there is a residual effect due to the extrinsic torsion. If that happens, then we will be able to compare the deduced value of $\kappa$ with observations of cosmic anisotropy in the matter content of the Universe, if these should be confirmed. The same quantity, $\varepsilon,$ describes these two apparently unrelated phenomena. In principle, then, the theory is falsifiable, in a rather remarkable manner.
We now turn to what is perhaps the most familiar form of cosmic anisotropy, that of the redshift field.
8. Anisotropy and the Redshift Field
While the redshift field is no longer the sole way of observing the Hubble field, it is still the principal way. We now wish to explore how anisotropies due to extrinsic torsion affect it. (Some relevant observational references are [64, 65]; see also the discussion of these results in [22].)
The situation here is quite different to the one we contemplated in the preceding Section. There, we were concerned with anisotropies in the “initial” data, with the idea that these might ultimately be observable indirectly, in the guise of their effects on the subsequent evolution. Here, however, we are interested in direct observations of photons propagating to us from very early times, when the torsion might have been significantly larger than it is now.
Before proceeding to that, however, we need a brief digression on the equation of a photon worldline in torsional theories. (The analysis in this digression is based on the relevant discussions in [66, 45].)
Let $u$ be the tangent vector to a null curve (usually called the four-dimensional wave vector), with a canonical parameter $\lambda,$ in a spacetime with torsion (not necessarily axial). In geometric optics, $u$ can be expressed as the four-dimensional gradient of a scalar eikonal function $\varphi$; that is, the one-form dual to $u$ (using the spacetime metric $g^{*}$), let us call it $U,$ can be written as $U=d\varphi.$ This is true not just for the standard Maxwell equations, but also for a very large family of generalisations of those equations [66].
The Hessian of $\varphi$, $\nabla^{*}d\varphi=\nabla^{*}U$, is not necessarily a symmetric bilinear form if the torsion is not zero: instead one has
$\nabla^{*}_{A}U(B)\,-\,\nabla^{*}_{B}U(A)\,=\,-\,U(T^{*}(A,\,B)),$ (42)
where $A$ and $B$ are arbitrary spacetime vectors [^17] With respect to a coordinate basis, we would write the components of $U$ as $u_{a}$, and so we see that this is simply the statement that $\nabla^{*}_{a}u_{b}-\nabla^{*}_{b}u_{a}=-T^{*c}_{ab}u_{c},$ where $u_{c}=g^{*}_{cd}u^{d}=\partial_{c}\varphi.$. (Note that the derivation of this equation relies on the fact that $U$ is an exact form; that is, on the assumption that $u$ is expressible in terms of an eikonal.)
Now set $A=u;$ then using the fact that the metric has zero covariant derivative we find that the second term on the left becomes $u\cdot\nabla^{*}_{B}u,$ and this is zero because $u\cdot u=0.$ (Here, for notational clarity, we are using the dot to represent the semi-Riemannian scalar product defined by $g^{*}.$) A similar calculation applied to the first term allows us finally to conclude that
$\nabla^{*}_{u}U\,=\,-\,U(T^{*}(u,\_)).$ (43)
Notice that the left side of this is just the dual one-form to $\nabla^{*}_{u}u.$ Relative to a coordinate basis, it is $u^{a}\nabla^{*}_{a}u_{b},$ and so equation (43) can be expressed in the perhaps rather more transparent form
$u^{a}\nabla^{*}_{a}u_{b}\,=\,-\,T^{*c}_{ab}u_{c}u^{a}$ (44)
with respect to a coordinate basis. This is the equation for the worldlines of photons. (It is also actually the equation for the worldline of a free massive particle.)
We see that these worldlines are not, in general, geodesics. This is in sharp contrast to the case of General Relativity, where photons, as free particles, are assumed to have worldlines which are “straight lines” in spacetime.
This is well known, but we have derived it here in order to emphasise that it can indeed be derived from the presumably uncontroversial assumptions that $u$ is null and is given by the gradient of an eikonal (and that $\nabla^{*}$ is compatible with $g^{*}$). We do not need to assume that we are using any specific torsional theory, such as the Einstein-Cartan theory; we do not even need to ask whether photons are sensitive to the existence of intrinsic spin, and so on.
Nevertheless it seems at first rather strange that torsional theories can apparently violate something as basic as Newton’s First Law —  which is essentially what we are saying here. But this is where the study of spacetime anisotropy comes to the fore. The intuition that free particles should move along straight lines is built on the assumption that all directions in space are “equally good”: a free particle cannot “prefer” one direction over another, so it has to continue along a straight path. That would have seemed natural to Newton, but, even in ordinary General Relativity, it is very questionable, since we know that three-dimensional space can be anisotropic in general —  and this anisotropy can play a fundamental role in the physics, as it does for example in the Schwarzschild geometry.
Thus, this version of the First Law rests on rather doubtful foundations even in ordinary General Relativity. In torsional theories, where, as we have seen, some kind of anisotropy is inescapable, we have even less reason to think that worldlines of free particles should be straight, so there is really no occasion for surprise when we find that they (often) are not [^18] Newton’s First Law can (and should) be salvaged by reformulating it: by declaring that a particle on which no external influence acts should have an action defined in terms of quantities pertaining only to the particle itself (its mass if any, the length or canonical parameter along its own worldline). The principle of stationary action then leads to the correct equation for the worldline, that is, to equation (44), provided that we accept that gravitation is not an “external influence”. In this way we can preserve the greatest conceptual achievement of General Relativity, the abolition of gravitational “force”..
We now turn to the question of the anisotropy of the redshift field in torsional theories (not just axial torsion theories).
Let us consider a null curve, with tangent field $u$, which is the worldline of a photon of frequency $\omega,$ as seen by the distinguished observers whose worldines are integral curves of $\xi.$ Then [67]
$\omega\,=\,-\,\xi\cdot u,$ (45)
where, as above, the dot denotes the scalar product defined by $g^{*}.$
We wish to understand how $\omega$ varies along the null curve. To this end, we compute $d\omega\over d\lambda$}, where $\lambda$ is the canonical affine parameter as above:
${d\omega\over d\lambda}\,=\,-\,\nabla^{*}_{u}(\xi\cdot u)\,=\,-\xi\cdot\nabla^{*}_{u}u\,-\,u\cdot\nabla^{*}_{u}\xi.$ (46)
As $u$ is a null vector, we can write it as $u=\omega\xi+\omega n,$ where $n$ is a unit spatial vector representing the direction in which the photon is moving in each spatial section through which it passes. (That is, it is a vector field along the worldline of the photon, a function of $\lambda$.) We think of its initial value as being chosen in such a manner that the photon eventually arrives at our location; so this initial vector, and therefore the subsequent values of $n$, depends on the direction in which we choose to look (though not necessarily in a simple way).
We begin with the second term on the right. Since (see the beginning of Section 3, above) $\nabla^{*}_{\xi}\xi=0,$ we have, using $u=\omega\xi+\omega n$ and the fact that $\xi$ is perpendicular to $\nabla_{n}\xi$,
$u\cdot\nabla^{*}_{u}\xi\,=\,-\,\omega^{2}n\cdot\nabla^{*}_{n}\xi\,=\,-\,\omega^{2}h(n,\,n),$ (47)
where $h$ is the second fundamental form, given, in the spatially isotropic case, by equation (25) above. Since $\varkappa$ is antisymmetric, we have simply $\omega^{2}h(n,\,n)=H\omega^{2},$ since $n$ is a unit spacelike vector. Of course, this term does not depend on $n$. (Incidentally, this use of spatial isotropy marks the only point in our entire argument where we use any kind of symmetry.)
On the other hand, the first term on the right in equation (46) can be simplified using equation (43) (in practice, by using equation (44)): after some calculation one finds that
$\xi\cdot\nabla^{*}_{u}u\,=\,-\,\omega^{2}\left(\beta(n,\,n)\,+\,b(n)\right),$ (48)
where $\beta$ and $b$ are the tangential and normal components of the “bulk torsion” we defined in equations (16) and (18), above.
Combining these results, we have
${d\omega\over d\lambda}\,=\,-\,\omega^{2}\left(H\,+\,\beta(n,\,n)\,+\,b(n)\right).$ (49)
Adopting our usual basis, with $\xi$ as one of the basis vectors, we can write this in a more explicit way as follows:
${d\omega\over d\lambda}\,=\,\omega^{2}\left(-\,H\,+\,T^{*}_{ij0}n^{i}n^{j}\,+\,T^{*}_{0i0}n^{i}\right),$ (50)
where $i$ and $j$ run from 1 to 3, where $n^{i}$ are the components of the unit spatial direction vector.
The first term on the right in equation (49) is the usual one in General Relativity; it is of course isotropic. The second and third terms are not, unless the torsion takes a special form (see below). If we regard (49) as a differential equation for $\omega,$ the final value of $\omega$ when the photon is observed will be found by evaluating a solution of that equation. In general, it will of course depend on the initial value of $n$. (In fact it will also depend on the subsequent values of $n$, see below.) If so, then the redshift field will be anisotropic. It is the bulk torsion which determines whether the redshift field is isotropic.
However, something unusual occurs when the torsion is axial: $\beta(n,\,n)+b(n)=-\,T^{*}_{ij0}n^{i}n^{j}-T^{*}_{0i0}n^{i}=0.$ Thus, despite the fact that neither the extrinsic nor the bulk torsion is necessarily zero here, the redshift field for axial torsion theories is isotropic. This is consistent with the fact that the worldlines of photons are geodesics in this situation, as one sees immediately from equation (44).
We have seen that axial torsion theories do give rise to another form of anisotropy, arising from the imprint axial torsion leaves on the “initial conditions” at reheating. Thus we see that it is possible for one component of the Hubble field (the redshift field in this case) to be completely isotropic, even though other components are not.
While this is an instructive subtlety, we ask the reader not to attach too much importance to it. We have been assuming that the torsion during the inflationary era, up to reheating, is axial; but, during reheating, the inflaton is converted to other forms of matter (but still leaves its imprint on those other forms). To be concrete, let us assume that the immediate post-inflationary matter can be represented as a fluid. As this fluid will inherit the intrinsic spin of the inflaton, it can be represented (very) approximately by a Weyssenhoff fluid (see for example [68]). The torsion corresponding (via the Einstein-Cartan field equations) to such a fluid is not axial.
In more detail: the spin tensor of a Weyssenhoff fluid takes the form $v\,\otimes\,\sigma,$ where $v$ is the unit tangent vector field to the worldlines of the fluid, and $\sigma$ is a two-form representing the spin, which is taken to reside in a spatial section defined by $v$ (meaning that the spacetime one-form $\sigma(v,\,\_)$ vanishes identically). We will assume, as above, that the initial state of the fluid is dictated by the state of the inflaton at the start of reheating: so it inherits the geometric structure we discussed above. That is, it will be cylindrically symmetric (as seen from a distinguished spatial section) around a distinguished axis. It is reasonable to surmise that this distinguished axis leaves its mark on the Weyssenhoff fluid in the form of the eigenvector of $\sigma$ that has zero eigenvalue. (This eigenvector must exist, for the same reason as in our discussion in Section 6.)
In the Einstein-Cartan context, the torsion is then
$T^{*}\,=\,8\pi G\,v\,\otimes\,\sigma.$ (51)
A key observation here is that there is no justification for taking $v=\xi;$ that is merely one possibility. Unfortunately it is difficult to predict what $v$ might be: for that, one would need a precise description of “torsional reheating”. This is beyond the scope of the present work.
In general, then, the fluid worldlines are not perpendicular to the distinguished spatial sections $\Sigma$ in which the Universe is homogeneous. One says that such a cosmology is “tilted” [69]. Tilted cosmologies have recently seen a revival: see [70] for a particularly interesting example, and [71] for stimulating observations on this question.
Let us write $v=v^{0}\xi+v^{i}e_{i},$ where the $e_{i}$ are basis vectors for tangent spaces to $\Sigma,$ $v^{0}\neq 0$ (since $v$ is timelike) and i=1, 2, 3i=1,\,2,\,3. Then the components of the bulk torsion are (using $\sigma(v,\,\_)=0$):
$b_{i}\,=\,-\,8\pi\,G\sigma_{ik}v^{k},$ (52)
where the $\sigma_{ij}$ are the components of $\sigma$ relative to the $e_{i},$ while
$\beta_{ij}\,=\,8\pi\,G\sigma_{ik}{v^{k}\over v^{0}}v_{j}.$ (53)
Thus we have
$\beta(n,\,n)+b(n)=8\pi\,G\,\sigma_{ik}v^{k}n^{i}\left({n^{j}v_{j}\over v^{0}}\,-\,1\right).$ (54)
The bracketed expression is a non-zero multiple of the energy of the photon as seen by the observers with unit tangent $v$, so it is not zero, and nor, in general [^19] Notice however the following possible exception to this analysis: perhaps the projection of $v$ to $\Sigma$ lies in the direction of the distinguished axis, that is, parallel to the eigenvector of $\sigma$ with zero eigenvalue. One then sees that $\beta(n,\,n)+b(n)$ is in fact zero in this case, so there is no fundamental anisotropy, just as in the case of axial torsion. We consider this unlikely, but we cannot rule it out entirely., is $\sigma_{ik}v^{k}n^{i}$, provided that not all of the $v^{i}$ are zero; so equation (57) indicates that there is an anisotropy in the redshift field if $v\neq\xi$.
We see that the redshift field can be anisotropic for a Weyssenhoff fluid precisely when the cosmology is tilted. We regard this as the normal case, and so, generically, the cosmology we are outlining here does predict some anisotropy in the redshift field, as well as in the overall Hubble field, in the immediate post-Inflation era.
To summarise, then: the redshift field is, in a torsional theory, almost certainly anisotropic after reheating. But the photons we observe do not emanate from matter during that period: they come from the later time of decoupling (or “recombination”). In the interim, the photons cannot stream freely, so it may well be that all signs of the immediate post-Inflationary anisotropy have been lost by the time decoupling occurs. So we have to ask: can any torsion which survives to the time immediately after decoupling give rise to still further anisotropy, which might be directly observable?
By the time of decoupling/recombination, when the CMB photons began their journey towards us, we can assume that the spacetime metric is approximately of the general FRW form, with Euclidean spatial sections and a scale factor $a(t)$, where $t$ is the usual proper cosmic time. Let us suppose that the torsion, in particular the bulk torsion, is not yet negligible at this time and for an interval afterwards. (We will not assume that it arises from a Weyssenhoff fluid at this relatively late time, when the fluid approximation has broken down; instead we return to considering torsion of an arbitrary form.)
The cosmological redshift is often explained in the following elementary way. When the photon arrives at our location, its wavelength has been “stretched” by the cosmic expansion, by a factor of $a_{\mathrm{final}}\over a_{\mathrm{initial}}.$ Therefore its frequency has been reduced by that same factor, and so we define a redshift factor $z$ in the usual way, $z\,=\,{\omega_{\mathrm{initial}}\over\omega_{\mathrm{final}}}-1\;=\;{a_{\mathrm{final}}\over a_{\mathrm{initial}}}-1.$ If this is so, if the final frequency depends only on the initial and final values of the scale factor, then it seems that no further anisotropy can arise.
But is this actually the case? Is it clear that whatever conditions the photon experiences, in the vast “bulk” of spacetime from decoupling to our location, have no effect on the redshift? We need to investigate [^20] The following argument is essentially the solution of the suitably generalised version of Problem 4 in Chapter 5 of [67]..
Using equation (49), we have
${d(a\omega)\over d\lambda}\,=\,\omega{da\over d\lambda}\,-\,a\omega^{2}\left(H\,+\,\beta(n,\,n)\,+\,b(n)\right).$ (55)
Since $H=(da/dt)/a$ in this case, and $\omega=dt/d\lambda,$ we have $a\omega H=da/d\lambda,$ so that there is a cancellation and we obtain
${d(a\omega)\over d\lambda}\,=\,-\,a\omega^{2}(\beta(n,\,n)\,+\,b(n)).$ (56)
It is interesting to note that, in deriving this, we did not actually use the fact that the spatial sections are Euclidean.
If we set both parts of the bulk torsion equal to zero, we obtain a remarkable result: $a\omega$ has the same value when the photon is received as it had initially. This validates the above elementary explanation of the redshift: at the point of observation, the wavelength of the photon has been “stretched” by a factor $a_{\mathrm{final}}/a_{\mathrm{initial}}$, so the frequency (and the energy) has been reduced accordingly, because $a\omega$ is constant along the worldline.
But what is the physics underlying the constancy of $a\omega?$ Let us try to explain this.
In a time-dependent spacetime geometry, there is no timelike Killing vector field, and consequently no law of conservation of energy. The photon simply loses energy at some points along its path, and gains it at others, depending on the way in which the geometry varies with time; and there is no reason to expect that the gains at one time will compensate for the losses at another.
If however we imagine for example a zero-torsion cosmology with a periodically fluctuating scale factor, then the fact that $a\omega$ is constant means that the photon always does regain, over a complete cycle, whatever energy it has lost: there is no “memory” of any losses or gains. This is the real reason for the fact that the final energy loss depends only on the ratio $a_{\mathrm{final}}/a_{\mathrm{initial}}$ in the usual argument. The absence of what might be called “cosmic hysteresis” in zero-torsion cosmology is the real physical meaning of the constancy of $a\omega.$
When the bulk torsion makes its presence felt, however, the situation is very different. When equation (56) is integrated, we obtain
${\omega_{\mathrm{final}}\over\omega_{\mathrm{initial}}}\,=\,{a_{\mathrm{initial}}\over a_{\mathrm{final}}}\,\exp\left(\,-\,\int^{\lambda_{\mathrm{final}}}_{\lambda_{\mathrm{initial}}}\omega\left(\beta(n,\,n)\,+\,b(n)\right)d\lambda\right)\,;$ (57)
so we find that the ratio of the final frequency to the initial frequency depends on the values taken by the bulk torsion, and on the values of the spatial direction field $n$ (and also on the frequency itself) all along the path —  not just on the initial and final values of the scale factor. When the bulk torsion is not zero, the final redshift can no longer be understood simply in terms of the stretching of wavelengths: the loss of energy depends on the geometry of the far-off and long-ago “bulk” of spacetime as the photon passes through it. In short, bulk torsion gives rise to “cosmic hysteresis”.
The dependence on $n$, and on the bulk torsion fields if (as is generically the case) they themselves are not isotropic, means that the redshift field in a torsional theory will be anisotropic in general.
Even if the bulk torsion is negligible currently and was so in the recent past, the hysteresis could still have a potentially detectable anisotropic effect on the final redshift if the bulk torsion was not negligible even for a brief period in the distant past. Specifically: if there was an appreciable bulk torsion at and shortly after decoupling, then equation (57) means that it might have had an observable effect on the redshift now, even if the bulk torsion has diluted, to the point of being unobservable, since that time.
The exact form taken by this anisotropy is difficult to predict, precisely because of the dependence on the values of $\beta,$ $b$, $\omega$, and $n$ along the worldline of the photon; the integral on the right side of equation (56) will be very difficult to evaluate in practice. No doubt a numerical investigation would be necessary to determine it. But the essential point is clear: in a torsional theory, the redshift field may well be slightly anisotropic.
We noted, in Section 1, that we proposed to leave to one side other, non-geometric, potential sources of cosmic anisotropy, including the possibility of bulk flows [34, 35]. We have seen, in our discussion of the Weyssenhoff fluid, that “tilt” arises very naturally in torsional theories, so the two kinds of anisotropy might well occur simultaneously and be intertwined. It might be challenging to separate the two effects.
9. Conclusion
We live in a time of cosmological uncertainty [15]. There are several ways to respond to this; among them, to search for alternatives to General Relativity.
These alternatives are often themselves classified according to the degree to which they are “radical”. We have argued [37] that the least radical alternative to General Relativity is to consider non-zero torsion : because this is hardly an “alternative” at all.
To see this very briefly, we remind the reader that the equation of geodesic deviation for an arbitrary connection $\nabla$ takes the form (see [27], Chapter VIII)
$\nabla_{u}\nabla_{u}D\;+\;\,\nabla_{u}\left[T(D,u)\right]\;+\;R(D,u)u\;=\;0,$ (58)
where $u$ is the tangent vector field to a family of geodesics, $D$ is the field with integral curves parametrising the family, and $T$ and $R$ are the torsion and curvature. It is obvious that torsion and curvature have equal status here, and there is no fundamental justification for neglecting either. Like spacetime curvature, spacetime torsion describes the geometry directly, without needing the mediation of field equations and coupling constants, as matter fields do. There are of course systems such that torsion is small and can be neglected, just as there are systems in which Newtonian gravitation is an adequate approximation to General Relativity. It would be risky to assume that Inflationary cosmology belongs to either category.
We have seen however that, if we introduce torsion into cosmology, then there is a price to be paid : in all but a few special cases, Inflation no longer completely isotropises the early Universe. But recent observations show that this may be precisely what we need —  if indeed one or more of the various kinds of cosmic anisotropy should be confirmed observationally, then torsion may be a way to reconcile such observations with Inflation. Not only that : if this works, torsion will clearly be by far the most natural way of effecting that reconciliation.
However, the operative word here is, as it should be, “if ”. We have seen that torsion can, extremely naturally, account for some forms of anisotropy, but by no means all. Anisotropy along more than one axis ; anisotropy that involves no rotation ; anisotropy such that the magnitude of the Hubble field is not minimal along the distinguished axis ; anisotropy not accompanied by anomalies in the early value of the Hubble parameter : any of these, if confirmed, might present a serious challenge to our proposed reconciliation of Inflation with anisotropy.

### v2

@@
 We put this intuition on a firm basis, by proving that if we [a] make use of an Inflationary theory in which Inflation isotropises spatial geometry —  this, of course, includes the vast majority of such theories —  and if [b] we insist on assuming that spacetime has a strictly metric geometry (one in which the geometry is completely determined by a metric tensor), then indeed all aspects of the “Hubble field” must be isotropic.
-Conversely, should a Hubble anisotropy be confirmed, then either we must contrive to build anisotropy into Inflation from the outset, or we will have to accept that spacetime geometry is not strictly metric.
-We argue that allowing spacetime torsion to be non-zero would be by far the most natural way to accommodate such observations.
-Such theories make firm predictions, as for example that there should be a correlation between the degree of anisotropy at the end of Inflation and a certain specific component of the Hubble tension.
+Conversely, should a Hubble anisotropy be confirmed, then either we must contrive to build anisotropy into Inflation (and into the geometry of space) from the outset, or we will have to accept that spacetime geometry is not strictly metric.
+We argue that the second option, implemented by allowing spacetime torsion to be non-zero, would be by far the most natural way to accommodate such observations.
+Such theories can reconcile non-isotropic matter distributions with a perfectly isotropic spatial geometry, and thus are able to reconcile Inflation with possibly observed anisotropies.
+They also allow us to reconcile the absence of anisotropy in one era (say, that of the CMB) with its presence in another.
 1. Inflation and the Hubble Field
@@
 However, Inflation is still called upon to explain (or at least render less mysterious) some much more basic observations.
-Among these is the strange fact that, while spacetime has an observably non-trivial geometry, it seems that, at cosmological scales, three-dimensional space does not.
-Much effort has been expended on searches for the curvature of three-dimensional space, to no avail thus far (see [4] for a clear discussion, and [5, 6] for the latest data on this).
+Among these is the strange fact that, while spacetime has an observably non-trivial geometry, it seems that, at cosmological scales, three-dimensional space does not [^1] The evidence for Euclidean spatial geometry is interpreted as such in the context of General Relativity.
+In this work, we propose to modify GR (a little), so, strictly speaking, we should not assert that there is strong evidence for Euclidean spatial geometry in the modified theory.
+However, it is not plausible that such a modified theory would conspire to mislead us, by somehow changing the CMB angular power spectrum, the BAO, and so on, in just such a manner as to cancel the effects of an actual non-trivial spatial geometry..
+Much effort has been expended on searches for the curvature of three-dimensional space, to no avail thus far; it is important to note that this result has now been confirmed by many different lines of investigation [4, 5], with unprecedented precision (see [6, 7] for clear general discussions and [8, 9, 10] for the latest data on this).
 We are entitled to ask why our spacetime has this extraordinary structure.
-And it is indeed extraordinary: the only way for the spatial sections to have an effectively Euclidean [^1] We use “Euclidean” to mean “devoid of non-trivial geometry”, avoiding the word “flat” because it has misleading connotations: we mean that the curvature is zero, but also that any other measure of geometric non-triviality in three-dimensional space, such as torsion, should vanish.
+And it is indeed extraordinary: the only way for the spatial sections to have an effectively Euclidean [^2] We use “Euclidean” to mean “devoid of non-trivial geometry”, (usually) avoiding the word “flat” because it has misleading connotations: we mean that the curvature is zero, but also that any other measure of geometric non-triviality in three-dimensional space, such as intrinsic torsion, should vanish.
 intrinsic geometry is if the spacetime curvature, evaluated on them, resides entirely in their extrinsic geometry.
@@
 It is then plausible that the enormous, de Sitter-like expansion characteristic of Inflation reduced the region of space we can observe to such a small subspace (see [11] for this) that it becomes indeed effectively Euclidean.
-That sets up just the right initial conditions for the part of spacetime we can observe to evolve to its present unusual state, with spatial sections dominated by extrinsic curvature.
+That sets up just the right initial conditions for the part of spacetime we can observe to evolve to its present state, with its unusual foliation.
 Notice that this argument eliminates all indications of non-trivial spatial geometry: it relies only on the assumption that the spatial sections have the structure of a differentiable manifold (and not, for example, that of a cone).
@@
 For example, it has sometimes been asked, in effect, whether Inflation really gives rise to enough expansion to “inflate away” the geometry of three-dimensional space: see for example [13] and particularly [14].
-From a completely different viewpoint, it has been pointed out that the field responsible for Inflation, the “inflaton”, does not itself “inflate away”; if, therefore, the inflaton is inherently anisotropic —  if, for example, it is a vector field —  then we might not expect Inflation to predict statistical isotropy (see the discussion at the end of [8]).
-However, these alternatives to the standard picture of Inflation, while interesting, seem rather contrived; in the case of “vector Inflation”, it is also difficult to arrange for the field to be stable.
+From a completely different viewpoint, it has been pointed out that the field responsible for Inflation, the “inflaton”, does not itself (necessarily) “inflate away”; if, therefore, the inflaton is inherently anisotropic —  if, for example, it is a vector field [15, 16] —  then we might not expect Inflation to predict statistical isotropy.
+(See the discussion at the end of [12]; see also [17] for some interesting recent developments in this field.)
+This concept of “vector Inflation” builds anisotropy into Inflation from the outset, and arranges for that anisotropy to survive the extreme inflationary expansion.
+But it is a somewhat radical departure from Inflation as usually conceived, and, more importantly, it means that we have to abandon the claim, discussed above, that Inflation explains the fact that our Universe appears to be foliated by Euclidean spacelike hypersurfaces.
 It seems to us that the peculiar spacetime structure we observe, with essentially all of the spacetime curvature (evaluated on spatial sections) concentrated in the extrinsic curvature of those sections, reflects the action of some profound physical process, and is no accident.
-We shall therefore, henceforth, cleave to the standard picture, as described above: we will take it that Inflation does indeed “inflate away” all non-trivial spatial geometry, so that the spatial sections are effectively exactly Euclidean, and therefore statistically isotropic.
-This approach also has the advantage of simplicity (and clarity).
-All this has recently become topical and relevant, because statistical isotropy has been cast into doubt by analyses of several data sets: see for example [11, 12, 13, 14].
-(See also Section 4.6 of [15] for a survey of these and other possible anomalies in the recent cosmic observations.) While this “statistical anisotropy” is by no means confirmed [16], recent observational advances promise to settle the question in the not very distant future.
+We shall therefore, henceforth, cleave to the standard picture, as described above: we will take it that Inflation does indeed “inflate away” all non-trivial spatial geometry, not approximately but essentially exactly, so that the spatial sections have an intrinsic geometry which is effectively exactly Euclidean, and therefore statistically isotropic.
+Our claim will be that, strange as it may seem, this picture can be reconciled with observations of cosmic anisotropy.
+All this has recently become topical and relevant, because statistical isotropy has in fact been cast into doubt by analyses of several data sets: see for example [18, 19, 20, 21].
+(See also Section 4.6 of [22] for a survey of these and other possible anomalies in the recent cosmic observations.) While this “statistical anisotropy” is by no means confirmed [23], it is difficult to believe that the agreement of all these diverse data is a “fluke” [24].
+In any case, recent observational advances promise to settle the question in the not very distant future.
 Since these new observations challenge the very foundations of the standard understanding of the early Universe, they certainly deserve attention.
@@
 In a similar spirit, studies of galaxy clustering and alignments have also been proposed [28] as probes of large-scale anisotropy.
-These remarkable new developments might well give us effective ways of studying statistical anisotropy, independently of the redshift field (which of course remains of great interest, and in which anisotropies may also exist, see for example [21] and especially the very clear presentation in [22]).
-(Thus the plural in the title of the present work.)
+These remarkable new developments might well give us effective ways of studying statistical anisotropy, independently of the redshift field (which of course remains of great interest, and in which anisotropies may also exist, see for example [29, 30, 31, 32, 33, 34, 35] and especially the very clear presentation in [36]).
+Thus the plural in the title of the present work.
 Instead of referring only to the redshift field, then, we shall speak of a Hubble field, the field representing the directed rate at which the galaxies actually separate as cosmic time passes.
@@
 In short, the intrinsic metric $g$ and the second fundamental form $h$ are related in various rather deep ways, both geometric and physical; so it is not out of the question that $h$ might inherit isotropy from $g$, so that the full set of “initial conditions” bequeathed by Inflation is indeed necessarily isotropic.
-To summarise: the fate of the Hubble field is connected to that of the intrinsic geometry, including its isotropisation, in rather subtle ways.
-We wish to argue from this that it is not obvious that Inflationary isotropisation of spatial sections necessarily entails an isotropic Hubble field.
 Our objective here is to understand this precisely.
 We will find that, in all conventional Inflationary theories (meaning those in which anisotropy is not built in from the outset), and in all theories of gravitation using strictly metric spacetime geometry (meaning geometry which is completely fixed by a semi-Riemannian metric and nothing else), the extrinsic geometry does necessarily share the symmetries of the intrinsic geometry: in all such theories, the Hubble field (and the matter fields) must isotropise if Inflation occurs.
-This is why it is so difficult to reconcile statistical anisotropy with Inflation.
+This is why it is so difficult to reconcile statistical anisotropy with (conventional) Inflation.
 However, we also show that this is really a consequence of the special relationship between intrinsic and extrinsic geometries in strictly metric spacetimes, and that it is not the case in the simplest and geometrically most natural extensions of GR, those in which the torsion is permitted to be non-zero [42, 43, 44].
 Our principal findings regarding cosmic isotropy in torsional Inflation are as follows.
-First, when torsion is present, it is indeed possible for intrinsic isotropy to co-exist with anisotropy of the Hubble field: in other words, torsion supplies a natural way of reconciling Inflation with (some) observations of statistical anisotropy.
+First, when torsion is present, it is indeed possible for a perfectly isotropic intrinsic spatial geometry to co-exist with anisotropy of the matter distribution and therefore of the Hubble field: in other words, torsion supplies a natural way of reconciling Inflation with (some) observations of statistical anisotropy.
+(The essential technical point here is that the symmetries of the matter field are correlated, via the constraint equations, with the extrinsic, not the intrinsic, geometry of the spatial sections.)
 Second, certain special kinds of torsion (the “axial” variety) make simple predictions as to the form taken by the Hubble field anisotropy: for example, in the axial case, there is always just one distinguished direction, and that direction is associated with a minimum of the magnitude of the Hubble field.
 Third, the extent of the anisotropy induced by torsion is measured by parameters which appear again in the generalised Friedmann equation.
-Thus for example axial torsion theories imply the existence of correlation between the Hubble anisotropy and a possible torsional component of the much-discussed “Hubble tension” [31, 32, 33]; this means that observations of statistical anisotropy might in principle allow (falsifiable) predictions as to the value of that component, and vice versa.
-Finally, torsion can and often does disrupt the familiar simple relation between cosmic redshift and the scale factor, leading in some cases to a very complex and subtle form of anisotropy in the redshift field.
-In summary, we argue that torsion offers the most natural path to reconciling Inflation with (some) possible observations of statistical anisotropy, while linking them to other cosmological anomalies, such as the Hubble tension.
+Thus for example axial torsion theories imply the existence of correlation between the Hubble anisotropy and a possible torsional component of the much-discussed “Hubble tension” [45, 46, 47, 48]; this means that observations of statistical anisotropy might in principle allow (falsifiable) predictions as to the value of that component, and vice versa.
+Finally, we will see that, in torsional theories, the level of anisotropy at one cosmic epoch (for example, as seen in the CMB) may not dictate the level observable at other times, potentially removing any conflict between data sets corresponding to differing eras.
 Before we begin, we should point out that we are concerned here exclusively with what might be called “geometric” anisotropies.
-One should bear in mind, however, the possibility of anisotropies due to other causes, such as those due to bulk flows [34, 35] or to the growth of non-linear structures [36].
+One should bear in mind, however, the possibility of anisotropies in the observations due to other causes, such as those due to bulk flows [49, 50] or to the growth of non-linear structures [51].
 These might well be just as important as geometric anisotropies, and they may be superimposed on the effects we discuss.
 We exclude them here only to avoid confusion.
-(See however Section 8, below, where we will see that this separation may not ultimately be fully tenable.)
 We begin with some simple observations regarding ordinary Inflation, in the context of standard General Relativity.
 2. What Inflation Does and Does Not
-To reiterate: the spatial sections of our Universe, those in which (apart from small peculiar velocities) the galaxies are distributed uniformly, appear (at present, and in the directly observable past) to be Euclidean, to a good approximation [4, 5, 6].
+To reiterate: the spatial sections of our Universe, those in which (apart from small peculiar velocities) the galaxies are distributed uniformly, appear (at present, and in the directly observable past) to be Euclidean, to a good approximation [6, 7, 8, 9, 10].
 And yet, cosmic spacetime is by no means close to being Minkowskian.
@@
 for vectors W,Z,X,YW,\,Z,\,X,\,Y tangential to $\Sigma$, and where $h$ is the (0, 2)(0,\,2) version of $\hat{h}$.
-Note (see [37]) that this expression for the extrinsic curvature is true for any spacetime geometry endowed with a spacetime metric and a compatible linear connection: it is valid whether or not the geometry is what we are calling “strictly metric”.
+Note (see [52]) that this expression for the extrinsic curvature (and also the Gauss formula relating it to the full spacetime curvature) is true for any spacetime geometry endowed with a spacetime metric and a compatible linear connection: it is valid whether or not the geometry is what we are calling “strictly metric”.
 (That is, it is valid even if $h$ is not a symmetric bilinear form; but in that case the ordering of the vectors in equation (9) must be carefully preserved.)
@@
 Instead, we want the geometry around $\gamma(t)$ also to be approximately isotropic for at least some interval, of non-zero length, of parameter values along $\gamma$.
-This means that the Killing vectors associated with the isotropic mappings must commute with $\mathcal{L}_{\xi}$.
+This means that the Lie derivatives with respect to the Killing vectors associated with the isotropic mappings must commute with $\mathcal{L}_{\xi}$.
 Therefore equation (10) implies that each isotropic mapping of $g$ is also an isotropic mapping of $h$, that is, of $\mathfrak{H},$ since the two are equal in the strictly metric case.
@@
 In the simplest cases one can easily confirm that $H$ is as expected.
-If we choose a (proper) time coordinate such that the unit normal is $\xi=\partial_{t},$ and set $g=a(t)^{2}\,f,$ where $a(t)$ is the scale factor and $f$ is a time-independent isotropic three-dimensional spatial metric tensor, then we have
+If we use Gaussian normal coordinates with a proper time coordinate $t$ (so that $\xi=\partial_{t}$ when both are evaluated on $\Sigma$) and set $g=a(t)^{2}\,f,$ where $a(t)$ is the scale factor and $f$ is a time-independent isotropic three-dimensional spatial metric tensor, then we have
 $\mathfrak{H}\,=\,{\dot{a}\over a}\,g,$ (12)
@@
 Notice also that, as expected, during Inflation all measures of the extrinsic geometry, including the Hubble tensor/parameter, increase as the intrinsic geometry approaches Euclidean geometry.
-Inflation of a strictly metric spacetime, then, is a process whereby the Hubble tensor grows (from a very small or zero initial value) to a tensor which, by the time reheating begins, has a large magnitude and is as isotropic as the spatial metric.
+Inflation in the case of a strictly metric spacetime, then, is a process whereby the Hubble tensor grows (from a very small or zero initial value) to a tensor which, by the time reheating begins, has a large magnitude and is as isotropic as the spatial metric.
 There are two remarkable aspects of this argument.
-The first is that we have used the word “Inflation” in its broadest possible sense: as referring to an early phase of explosive expansion.
-The argument therefore applies to all of the many conventional (scalar) Inflationary models, and so the conclusion cannot be avoided by tinkering with the details (the precise nature of the inflaton potential or other features of its Lagrangian, and so on).
+The first is that we have used the word “Inflation” in its broadest possible sense: as referring to an early phase of extreme expansion.
+The argument therefore applies to all of the many conventional Inflationary models, and so the conclusion cannot be avoided by tinkering with the details (the precise nature of the inflaton potential or other features of its Lagrangian, and so on).
 Second, and still more striking: the remainder of the argument is purely geometric —  it makes no reference whatever to the field equations for gravity.
@@
 [b] The geometry of spacetime is not strictly metric.
-We wish to argue that option [a] should be rejected: the successes of conventional Inflation are too numerous and various.
-Therefore option [b] would be necessary if statistical anisotropy were to be confirmed observationally.
+We wish to argue that pursuing option [a] should be attempted only if all else fails.
+If we accept this, then option [b] would be necessary if statistical anisotropy were to be confirmed observationally.
 The question is whether it is sufficient: does a non-strictly metric spacetime geometry permit a statistical anisotropy of some kind, even if Inflation has isotropised the intrinsic spatial geometry?
@@
 Whether it can do so in a manner consistent with (all of) the observations is of course another matter.
-In fact, we will see that option [b] does permit some kinds of anisotropy, but, at least in the concrete examples we will study, not all.
+In fact, we will see that option [b] does permit some kinds of anisotropy, but not all.
 5. Isotropisation and Torsion: General Theory
@@
 The $\eta$ tensor is constructed from this, as usual, according to equation (19).
-Like $h$, $\mathfrak{H}$ is not, generically [^12] One can contrive a situation in which the antisymmetric part of $\beta$ exactly cancels $\varkappa$, so that $\mathfrak{H}$ becomes symmetric.
-Clearly this is a very special, artificial arrangement, and the reader can readily verify that it cannot happen in the cases of physical interest we will study below., a symmetric bilinear form when the torsion is not zero.
+Like $h$, $\mathfrak{H}$ is not, generically [^13] One can however force it to be symmetric by imposing isotropy “by hand”: see below., a symmetric bilinear form when the torsion is not zero.
 We saw earlier (Section 3, above) that this means that a non-trivial rotational component, $Q$ in equation (6), must be present.
@@
 For otherwise we would have to explain why $h$ takes some specific form at the beginning, and this mean importing a large number of seemingly arbitrary parameters into the theory.
-The same reasoning applies here, and so, if we accept this line of argument, then both terms on the right of equation (22) must (independently) vanish at that time.
+The same reasoning applies here, and so, if we accept this line of argument, then both terms on the right of equation (22) must (independently) vanish at the beginning.
 Both terms will however grow as Inflation proceeds, so, in particular, the extrinsic torsion $\varkappa$ will not be negligible by the time of reheating: as we will see in detail later, it must grow, to compensate for the loss of the intrinsic torsion.
@@
 In other words, the extrinsic geometry is automatically anisotropic if the extrinsic torsion does not vanish.
-Similarly, we have no reason to think that $\beta$ will isotropise; it certainly does not do so in the concrete examples we will consider [^13] One can however ask what happens if we force a torsional geometry to be highly symmetric.
-This question has led to important classification theorems: see [52, 53, 54, 55].
-Our point here, however, is just that Inflation alone does not impose such a high degree of symmetry on spacetime..
+(A similar argument implies that the intrinsic torsion has to vanish if it is to be isotropic.)
+Similarly, we have no reason to think that $\beta$ will isotropise; it certainly does not do so in the concrete examples we will consider.
+If we force it to be isotropic, then, as we have been discussing, its antisymmetric part has to vanish, and Schur’s lemma implies that its symmetric part is a time-dependent multiple, let us call it $4\phi(t)$, of $g$.
+If both the extrinsic and the “bulk” torsion are forced to be isotropic, then, both $h$ and $\mathfrak{H}$ are symmetric bilinear forms.
+(Note that the “bulk” torsion one-form $b$, given by equation (18), must vanish in the isotropic case because it is metrically dual to a vector field.)
+This special case of “isotropic torsion”, in which all forms of torsion vanish except for the “bulk” torsion bilinear form $\beta,$ given here by $\beta=4\phi g,$ is the case usually assumed in discussions of torsional cosmology: see [67].
+The question as to what happens if we force spacetime torsion to share the symmetries of the metric has been studied in a more general and complete manner, leading to important classification theorems: see [68, 69, 70, 71].
+Our point here, however, is just that Inflation alone does not impose such high degrees of symmetry on spacetime torsion.
 In short: the Hubble field is always isotropic in a zero-torsion Inflationary theory which isotropises the spatial metric; but (unless we force it to be so, “by hand”) it is never isotropic in an inflationary theory with non-vanishing torsion.
@@
 First, the spatial sections will, if $\varkappa\neq 0,$ embed into spacetime in an anisotropic way, specified by the detailed structure of $\varkappa$.
-Secondly, if $\beta\neq 0,$ anisotropies in the “bulk” (regions of spacetime remote from us) can give rise to anisotropic effects in signals we observe, depending on the properties of $\beta$.
+Secondly, if $\beta\neq 0,$ anisotropies in the “bulk” (regions of spacetime remote from us) can and normally will give rise to anisotropies, which (as we will see later) can appear in signals we observe, arriving from afar.
 This is why it is natural to propose that torsion may help us to reconcile Inflation with purported statistical anisotropies of various kinds: torsion provides us with a large (though, as we saw, not limitless) repertoire of techniques to induce anisotropy in those parts of the Hubble field not controlled by the intrinsic geometry of spatial sections.
@@
 $\mathfrak{H}\,=\,Hg\,+\,{1\over 2}\,\varkappa\,+\,{1\over 2}\,\beta.$ (26)
+In the exceptional case of “isotropic torsion”, this equation takes the form $\mathfrak{H}=\left(H\,+\,2\phi\right)g$.
+(The coefficient of $\phi$ was chosen so as to ensure agreement with [67].) Thus “isotropic torsion” makes its presence felt through a modification of the Hubble parameter.
+We will see later that a similar effect arises in the anisotropic case.
 The consequences of having an anisotropic Hubble tensor is most clearly explained if we focus on torsion (and contortion) tensors of a special kind.
@@
 $Q]\,=\,\left(\begin{array}[]{ccc}{H\over\sqrt{H^{2}+\kappa^{2}/4}}&{-\kappa/2\over\sqrt{H^{2}+\kappa^{2}/4}}&0\\ {\kappa/2\over\sqrt{H^{2}+\kappa^{2}/4}}&{H\over\sqrt{H^{2}+\kappa^{2}/4}}&0\\ 0&0&1\\ \end{array}\right).$ (32)
-As we foresaw, this rotation is always non-trivial if the extrinsic torsion does not vanish. This “Hubble twist” is characteristic of the presence of extrinsic torsion.
+As we foresaw, this rotation is always non-trivial if the extrinsic torsion does not vanish. This “Hubble twist” is particularly simple in the case of axial torsion.
 The rotation is clearly around an axis which coincides with the distinguished axis of the $\eta$ tensor.
@@
 $\varepsilon\,=\,{\kappa\over H};$ (33)
-This gives us a dimensionless measure of the anisotropy; it is also a measure of the relative importance of extrinsic torsion to extrinsic curvature.
+This gives us a dimensionless measure of the anisotropy; it is also a measure of the relative importance of extrinsic torsion compared to extrinsic curvature.
 In terms of $\varepsilon,$ the angle of the “Hubble twist” is given by
@@
 But it is instructive, in the sense that it clarifies the kinds of effects to which torsion can give rise.
-We can see four facts very plainly, in the case of axial torsional theories of cosmic anisotropy.
+We can see four facts very plainly, in the case of axial torsional theories of the geometry of spacetime during Inflation.
 [a] The predicted anisotropy singles out a distinguished direction; the Hubble parameter $H$ is the rate of separation in that direction.
 The rates of separation in other directions are still controlled by $H,$ so we can say that the Hubble parameter is still a measure of the overall rate of separation of the galaxies.
-It coincides with the “Hubble parameter” in discussions where the anisotropy is neglected, as for example in the study of the “Hubble tension”.
+It plays the role of the “Hubble parameter” in discussions where the anisotropy is neglected, as for example in the study of the “Hubble tension”.
 [b] We saw earlier that not every form of anisotropy can be explained by extrinsic torsion, because torsional anisotropy always has a rotational component.
@@
 This, then, is the predicted form of the geometric anisotropy at the end of Inflation.
-This geometric anisotropy would have left traces on the matter distribution at reheating.
+This geometric anisotropy might have left traces on the matter distribution at reheating.
 We now ask exactly how that works; the hope being that, by working backwards from current observations, one will obtain a theory that can be falsified by deductions potentially in conflict with [a], [b], [c], or [d].
+(It is important to notice, however, that some observations of cosmic anisotropy are concerned with spacetime geometry long after the end of Inflation, so some of these remarks may not apply to them.
+This is particularly the case for the redshift field, as we will discuss below in Section 8.)
 7. How Torsional Inflation Sets up Initial Conditions for Reheating
@@
 here, as before, $\rho$ is the energy density, but now $S$ is the scalar curvature of the full linear connection, including spatial torsion and its derivatives, on the spatial section.
-Note that, in the presence of (extrinsic) torsion, $h$ is not a symmetric bilinear form, so the final term is not the square of $\hat{h}$; this proves to be important (see [37]).
+Note that, in the presence of (extrinsic) torsion, $h$ is not a symmetric bilinear form, so $\hat{h}^{T}\hat{h}$ is not the square of $\hat{h}$; this proves to be important (see [52]).
 Since we are assuming that all geometric quantities on the spatial section have inflated away by the end of Inflation, we set $S=0$.
@@
 Equation (40) therefore tells us that the energy density at the time when reheating is about to start is also not isotropic.
-The anisotropy of the extrinsic geometry leaves its mark on the “initial” energy density, even though the underlying space is isotropic; and this mark may persist to recent times, as the system evolves.
+The anisotropy of the extrinsic geometry leaves its mark on the “initial” energy density, even though the underlying space is isotropic; and this mark may persist to later times, as the system evolves.
 One might try, for example, to use this general idea to develop a theory of the initial conditions leading to the anisotropies which may have been observed in massive galactic haloes [25, 26, 27, 28], independently of whatever one observes in the redshift field (which we about to discuss in the next Section).
@@
 8. Anisotropy and the Redshift Field
-While the redshift field is no longer the sole way of observing the Hubble field, it is still the principal way.
-We now wish to explore how anisotropies due to extrinsic torsion affect it.
-(Some relevant observational references are [64, 65]; see also the discussion of these results in [22].)
+While the redshift field is no longer the sole way of observing the Hubble field, it is still the principal way. We now wish to explore how anisotropies due to torsion affect it.
 The situation here is quite different to the one we contemplated in the preceding Section.
 There, we were concerned with anisotropies in the “initial” data, with the idea that these might ultimately be observable indirectly, in the guise of their effects on the subsequent evolution.
-Here, however, we are interested in direct observations of photons propagating to us from very early times, when the torsion might have been significantly larger than it is now.
+Here, however, we are interested in direct observations of photons propagating to us from various early times, when the torsion might have been significantly larger than it is now.
 Before proceeding to that, however, we need a brief digression on the equation of a photon worldline in torsional theories.
 (The analysis in this digression is based on the relevant discussions in [80, 60].)
-Let $u$ be the tangent vector to a null curve (usually called the four-dimensional wave vector), with a canonical parameter $\lambda,$ in a spacetime with torsion (not necessarily axial).
+Let $u$ be the tangent vector, usually called the four-dimensional wave vector, to a null curve with a parameter $\lambda$ (which we can modify if necessary) in a spacetime with torsion (of any kind).
 In geometric optics, $u$ can be expressed as the four-dimensional gradient of a scalar eikonal function $\varphi$; that is, the one-form dual to $u$ (using the spacetime metric $g^{*}$), let us call it $U,$ can be written as $U=d\varphi.$ This is true not just for the standard Maxwell equations, but also for a very large family of generalisations of those equations [80].
@@
 $u^{a}\nabla^{*}_{a}u_{b}\,=\,-\,T^{*c}_{ab}u_{c}u^{a}$ (44)
-with respect to a coordinate basis. This is the equation for the worldlines of photons. (It is also actually the equation for the worldline of a free massive particle.)
-We see that these worldlines are not, in general, geodesics.
-This is in sharp contrast to the case of General Relativity, where photons, as free particles, are assumed to have worldlines which are “straight lines” in spacetime.
+with respect to a coordinate basis, where of course the coefficients $T^{*c}_{ab}$ are the components of the spacetime torsion tensor.
+This is the equation for the worldlines of photons.
+(It is also actually the equation for the worldline of a free massive particle.)
+We see that these worldlines are not, in general, geodesics (though they can be).
+This is in sharp contrast to the case of General Relativity, where photons, as free particles, are assumed to have worldlines which are always “straight lines” in spacetime.
 This is well known, but we have derived it here in order to emphasise that it can indeed be derived from the presumably uncontroversial assumptions that $u$ is null and is given by the gradient of an eikonal (and that $\nabla^{*}$ is compatible with $g^{*}$).
 We do not need to assume that we are using any specific torsional theory, such as the Einstein-Cartan theory; we do not even need to ask whether photons are sensitive to the existence of intrinsic spin, and so on.
-Nevertheless it seems at first rather strange that torsional theories can apparently violate something as basic as Newton’s First Law —  which is essentially what we are saying here.
-But this is where the study of spacetime anisotropy comes to the fore.
-The intuition that free particles should move along straight lines is built on the assumption that all directions in space are “equally good”: a free particle cannot “prefer” one direction over another, so it has to continue along a straight path.
-That would have seemed natural to Newton, but, even in ordinary General Relativity, it is very questionable, since we know that three-dimensional space can be anisotropic in general —  and this anisotropy can play a fundamental role in the physics, as it does for example in the Schwarzschild geometry.
-Thus, this version of the First Law rests on rather doubtful foundations even in ordinary General Relativity.
-In torsional theories, where, as we have seen, some kind of anisotropy is inescapable, we have even less reason to think that worldlines of free particles should be straight, so there is really no occasion for surprise when we find that they (often) are not [^18] Newton’s First Law can (and should) be salvaged by reformulating it: by declaring that a particle on which no external influence acts should have an action defined in terms of quantities pertaining only to the particle itself (its mass if any, the length or canonical parameter along its own worldline).
+Nevertheless it seems at first rather strange that torsional theories can apparently violate something as basic as Newton’s First Law (“A free particle moves along a straight line at constant speed.”) —  which is essentially what we are saying here.
+But this is where the study of anisotropy comes to the fore.
+The intuition that free particles should have straight worldlines is built on the assumption that space and time are (in a sense that can readily be made precise) uniform: a free particle cannot in that case “justify” deviating from straight line motion, or moving at a non-constant speed.
+This assumption of uniformity seemed natural to Newton, and it also holds true in Special Relativity; but, even in ordinary General Relativity, it does not.
+For we know that, in that theory, three-dimensional space is not “uniform” in general —  and this can play a fundamental role in the physics, as it does for example in the Schwarzschild geometry.
+Similarly, time is not “uniform” in spacetimes, such as cosmological spacetimes, that lack a timelike Killing vector field.
+Thus, this version of the First Law rests on very doubtful foundations even in ordinary General Relativity.
+In torsional theories, where, as we have seen, some kind of spatial anisotropy is almost inescapable, we have even less reason to think that worldlines of free particles should be straight, so there is really no occasion for surprise when we find that they (often) are not [^18] Newton’s First Law can (and should) be salvaged by reformulating it: by declaring that a particle on which no external influence acts should have an action defined in terms of quantities pertaining only to the particle itself (its mass if any, the length or canonical parameter along its own worldline).
 The principle of stationary action then leads to the correct equation for the worldline, that is, to equation (44), provided that we accept that gravitation is not an “external influence”.
 In this way we can preserve the greatest conceptual achievement of General Relativity, the abolition of gravitational “force”..
-We now turn to the question of the anisotropy of the redshift field in torsional theories (not just axial torsion theories).
-Let us consider a null curve, with tangent field $u$, which is the worldline of a photon of frequency $\omega,$ as seen by the distinguished observers whose worldines are integral curves of $\xi.$ Then [67]
+In the special case of “isotropic torsion”, the right side of equation (44) is $(\phi\xi^{a}u_{a})u_{b}$ (see [67]), which is of the form (function of the parameter $\lambda$) $\times\,u_{b}$.
+In this case, it is well known that a change of parameter reduces the right side of (44) to zero, so the worldline of a photon in the isotropic case is in fact a geodesic.
+This underlines the claim that the worldlines of photons are usually not geodesics in torsional theories, because such theories usually involve some kind of anisotropy [^19] The precise statement is that when a torsional spacetime is isotropic, then the worldlines of particles are geodesics.
+The converse is not true, because sometimes anisotropy is present but the particles may not be able to detect it.
+This happens, in fact, when the torsion is axial: see below..
+We now turn to the question of the anisotropy of the redshift field in torsional theories.
+Let us consider a null curve, with tangent field $u$, which is the worldline of a photon of frequency $\omega,$ as seen by the distinguished observers whose worldines are integral curves of $\xi.$ As $u$ is a null vector, and because the timelike component of $u$ is the frequency [81], we can write it as $u=\omega\xi+\omega n,$ where $n$ is a unit spatial vector representing the direction in which the photon is moving in each spatial section through which it passes.
+(That is, it is a vector field along the worldline of the photon, a function of $\lambda$, just as $\omega$ is a scalar function of $\lambda$.) We think of its initial value as being chosen in such a manner that the photon eventually arrives at our location; so this initial vector can be regarded as dependent on the direction in which we choose to look (though not necessarily in a simple way).
+Equation (43) (or (44)) is a differential equation for $u$, but we can think of it as a pair of simultaneous ordinary differential equations for $\omega$ and $n$, as functions of $\lambda$.
+Normally the variables in such a pair of non-linear, intertwined equations cannot be separated; if that is so, the final value of $\omega$ depends not only on the initial value of $\omega$ but also on the initial value of $n$.
+In that case, the final, observed value of $\omega$ depends on the direction in which we choose to look: in other words, the redshift field is anisotropic.
+Let us see how this works in detail. First, we must find the equation for $\omega$ explicitly.
+We have
 $\omega\,=\,-\,\xi\cdot u,$ (45)
 where, as above, the dot denotes the scalar product defined by $g^{*}.$
-We wish to understand how $\omega$ varies along the null curve. To this end, we compute $d\omega\over d\lambda$}, where $\lambda$ is the canonical affine parameter as above:
+We wish to compute $d\omega\over d\lambda$}, where $\lambda$ is the canonical affine parameter as above:
 ${d\omega\over d\lambda}\,=\,-\,\nabla^{*}_{u}(\xi\cdot u)\,=\,-\xi\cdot\nabla^{*}_{u}u\,-\,u\cdot\nabla^{*}_{u}\xi.$ (46)
-As $u$ is a null vector, we can write it as $u=\omega\xi+\omega n,$ where $n$ is a unit spatial vector representing the direction in which the photon is moving in each spatial section through which it passes.
-(That is, it is a vector field along the worldline of the photon, a function of $\lambda$.) We think of its initial value as being chosen in such a manner that the photon eventually arrives at our location; so this initial vector, and therefore the subsequent values of $n$, depends on the direction in which we choose to look (though not necessarily in a simple way).
 We begin with the second term on the right.
@@
 $u\cdot\nabla^{*}_{u}\xi\,=\,-\,\omega^{2}n\cdot\nabla^{*}_{n}\xi\,=\,-\,\omega^{2}h(n,\,n),$ (47)
-where $h$ is the second fundamental form, given, in the spatially isotropic case, by equation (25) above.
-Since $\varkappa$ is antisymmetric, we have simply $\omega^{2}h(n,\,n)=H\omega^{2},$ since $n$ is a unit spacelike vector.
-Of course, this term does not depend on $n$.
-(Incidentally, this use of spatial isotropy marks the only point in our entire argument where we use any kind of symmetry.)
+where $h$ is the second fundamental form. (The appearance of $h$ here is how the Hubble field will enter the discussion.)
 On the other hand, the first term on the right in equation (46) can be simplified using equation (43) (in practice, by using equation (44)): after some calculation one finds that
 $\xi\cdot\nabla^{*}_{u}u\,=\,-\,\omega^{2}\left(\beta(n,\,n)\,+\,b(n)\right),$ (48)
-where $\beta$ and $b$ are the tangential and normal components of the “bulk torsion” we defined in equations (16) and (18), above.
-Combining these results, we have
-${d\omega\over d\lambda}\,=\,-\,\omega^{2}\left(H\,+\,\beta(n,\,n)\,+\,b(n)\right).$ (49)
+where $\beta$ and $b$ are the tangential and normal components of the “bulk” torsion we defined in equations (16) and (18), above.
+Combining these results, we have (simplifying by using $\omega^{-1}$ instead of $\omega$)
+${d\omega^{-1}\over d\lambda}\,=\,h(n,\,n)\,+\,\beta(n,\,n)\,+\,b(n).$ (49)
 Adopting our usual basis, with $\xi$ as one of the basis vectors, we can write this in a more explicit way as follows:
-${d\omega\over d\lambda}\,=\,\omega^{2}\left(-\,H\,+\,T^{*}_{ij0}n^{i}n^{j}\,+\,T^{*}_{0i0}n^{i}\right),$ (50)
+${d\omega^{-1}\over d\lambda}\,=\,h_{ij}n^{i}n^{j}\,+\,T^{*}_{i0j}n^{i}n^{j}\,+\,T^{*}_{00i}n^{i},$ (50)
 where $i$ and $j$ run from 1 to 3, where $n^{i}$ are the components of the unit spatial direction vector.
-The first term on the right in equation (49) is the usual one in General Relativity; it is of course isotropic.
-The second and third terms are not, unless the torsion takes a special form (see below).
-If we regard (49) as a differential equation for $\omega,$ the final value of $\omega$ when the photon is observed will be found by evaluating a solution of that equation.
-In general, it will of course depend on the initial value of $n$.
-(In fact it will also depend on the subsequent values of $n$, see below.) If so, then the redshift field will be anisotropic.
-It is the bulk torsion which determines whether the redshift field is isotropic.
-However, something unusual occurs when the torsion is axial: $\beta(n,\,n)+b(n)=-\,T^{*}_{ij0}n^{i}n^{j}-T^{*}_{0i0}n^{i}=0.$ Thus, despite the fact that neither the extrinsic nor the bulk torsion is necessarily zero here, the redshift field for axial torsion theories is isotropic.
-This is consistent with the fact that the worldlines of photons are geodesics in this situation, as one sees immediately from equation (44).
+Clearly, in general, $n$ is very much present on the right side of equation (49), and $\omega$ cannot be separated from it; and so, integrating this equation, we will find that the final value of $\omega$, that is, the redshift field, is generically anisotropic, especially if “bulk” torsion is present.
+In particular, $b$, being essentially a spatial vector field, cannot be isotropic if it is not zero.
+We have said that anisotropy corresponds to a dependence of the final value of $\omega$ on the “initial” value of $n$.
+But $n$ is not a constant along the worldline in general, so this “initial” value will depend on our choice as to when the “initial” time should be.
+This means that the extent of the anisotropy in the redshift field can and usually will be time-dependent in the presence of torsion.
+We will return to this important observation at the end of this Section.
+If we restrict to a strictly metric theory, such as General Relativity, then the torsion terms on the right side of equation (49) are zero.
+If we assume that Inflation performs its customary service of isotropising the spatial sections, then, as usual, the second fundamental form inherits isotropy, and this is the essential point: we have $h=Hg$, and so, since $n$ is a unit vector, this gives us h(n,n)=H,h(n,\,n)=H, where now $H$ is to be regarded as a function of $\lambda$ (only).
+Then (49) is greatly simplified, and when integrated it gives
+$\omega^{-1}\,=\,\int\,H\,d\lambda.$ (51)
+All dependence on $n$ has dropped out; the change in the value of $\omega$ is due exclusively to the cumulative effect of the “stretching” described by the Hubble parameter, which is inherently isotropic.
+It follows that, in conventional Inflationary theories embedded in General Relativity or another strictly metric theory, the redshift field inherits isotropy from the spatial geometry, in the familiar manner.
+(A similar argument works in the case of “isotropic torsion” [67]: again, the redshift field is always isotropic in that theory if the intrinsic spatial geometry is so, since $b(n)=0$ and $\beta(n,\,n)=4\phi g,$ so the integrand in equation (49) is independent of $n$: it is equal to $H\,+\,2\phi,$ and equation (51) is valid if modified in the same way.)
+Thus, in strictly metric theories of gravitation, redshift anisotropy can only be present if it occurs in the earliest relevant times, that is, in the CMB; but, again, that would be hard to reconcile with Inflation.
+Furthermore, such anisotropies would tend to dissipate with the passing of time [12] (but see the potentially very important novel viewpoint on this put forward in [82]).
+This is presumably why, when, in [29], substantial anisotropies were reported, it was claimed that these observations had “already been refuted” by the fact that anisotropies at that level (or higher) were not seen in the CMB.
+We will argue that this refutation was indeed correct, but only if torsion is neglected.
+Returning to the torsional case, we have, when the intrinsic geometry of the spatial sections is isotropic,
+$\omega^{-1}\,=\,\int\,\left(H\,+\,\beta(n,\,n)\,+\,b(n)\right)\,d\lambda.$ (52)
+We stress again that the right side [^20] We notice in passing that the extrinsic torsion plays no role here : it drops out because the extrinsic torsion form is antisymmetric (see equation (22)).
+So it is still the case that h(n,n)=H.h(n,\,n)=H.
+Instead the kind of torsion responsible for redshift anisotropy is the “bulk” torsion.
+tells us both that the redshift field is generically anisotropic in the presence of torsion, and that the “anisotropy field” varies with time.
+However, something unusual occurs when the torsion is axial : $\beta(n,\,n)+b(n)=T^{*}_{i0j}n^{i}n^{j}\,+\,T^{*}_{00i}n^{i}=0:$ torsion is present, but photons are insensitive to it.
+Thus, despite the fact that neither the extrinsic nor the bulk torsion is necessarily zero here, the redshift field for axial torsion theories is isotropic in a given epoch if the intrinsic spatial geometry is isotropic then.
 We have seen that axial torsion theories do give rise to another form of anisotropy, arising from the imprint axial torsion leaves on the “initial conditions” at reheating.
@@
 To be concrete, let us assume that the immediate post-inflationary matter can be represented as a fluid.
-As this fluid will inherit the intrinsic spin of the inflaton, it can be represented (very) approximately by a Weyssenhoff fluid (see for example [68]).
+As this fluid will inherit the intrinsic spin of the inflaton, it can be represented approximately by a Weyssenhoff fluid (see for example [83]).
 The torsion corresponding (via the Einstein-Cartan field equations) to such a fluid is not axial.
 In more detail : the spin tensor of a Weyssenhoff fluid takes the form $v\,\otimes\,\sigma,$ where $v$ is the unit tangent vector field to the worldlines of the fluid, and $\sigma$ is a two-form representing the spin, which is taken to reside in a spatial section defined by $v$ (meaning that the spacetime one-form $\sigma(v,\,\_)$ vanishes identically).
-We will assume, as above, that the initial state of the fluid is dictated by the state of the inflaton at the start of reheating: so it inherits the geometric structure we discussed above.
-That is, it will be cylindrically symmetric (as seen from a distinguished spatial section) around a distinguished axis.
-It is reasonable to surmise that this distinguished axis leaves its mark on the Weyssenhoff fluid in the form of the eigenvector of $\sigma$ that has zero eigenvalue.
-(This eigenvector must exist, for the same reason as in our discussion in Section 6.)
 In the Einstein-Cartan context, the torsion is then
 $T^{*}\,=\,8\pi G\,v\,\otimes\,\sigma.$ (53)
+We will assume, as above, that the initial state of the fluid is dictated by the state of the inflaton at the start of reheating : so it inherits the geometric structure we discussed above.
+In fact, $\sigma$ is just a non-zero multiple of the extrinsic torsion form at this time, and it follows that its non-zero eigenvalue will be cylindrically symmetric (as seen from a distinguished spatial section) around a distinguished axis, defined by the eigenvector of $\sigma$ that has zero eigenvalue.
 A key observation here is that there is no justification for taking $v=\xi;$ that is merely one possibility.
@@
 One says that such a cosmology is “tilted” [84].
-Tilted cosmologies have recently seen a revival: see [70] for a particularly interesting example, and [71] for stimulating observations on this question.
+Tilted cosmologies have recently seen a revival : see [82] for a particularly interesting example, and [85] for stimulating observations on related questions.
 Let us write $v=v^{0}\xi+v^{i}e_{i},$ where the $e_{i}$ are basis vectors for tangent spaces to $\Sigma,$ $v^{0}\neq 0$ (since $v$ is timelike) and i=1, 2, 3i=1,\,2,\,3.
@@
 We regard this as the normal case, and so, generically, the cosmology we are outlining here does predict some anisotropy in the redshift field, as well as in the overall Hubble field, in the immediate post-Inflation era.
+In the context of standard General Relativity, tilted cosmologies can account for anisotropies in the observations [82], but this is achieved at the expense of having anisotropic spatial sections; as we know, this cannot be avoided in that case.
+The advantage of the torsional approach is that one can obtain the same objective without spatial anisotropy, in accord with Inflation.
 To summarise, then : the redshift field is, in a torsional theory, almost certainly anisotropic after reheating.
-But the photons we observe do not emanate from matter during that period: they come from the later time of decoupling (or “recombination”).
-In the interim, the photons cannot stream freely, so it may well be that all signs of the immediate post-Inflationary anisotropy have been lost by the time decoupling occurs.
-So we have to ask: can any torsion which survives to the time immediately after decoupling give rise to still further anisotropy, which might be directly observable?
-By the time of decoupling/recombination, when the CMB photons began their journey towards us, we can assume that the spacetime metric is approximately of the general FRW form, with Euclidean spatial sections and a scale factor $a(t)$, where $t$ is the usual proper cosmic time.
-Let us suppose that the torsion, in particular the bulk torsion, is not yet negligible at this time and for an interval afterwards.
-(We will not assume that it arises from a Weyssenhoff fluid at this relatively late time, when the fluid approximation has broken down; instead we return to considering torsion of an arbitrary form.)
-The cosmological redshift is often explained in the following elementary way.
-When the photon arrives at our location, its wavelength has been “stretched” by the cosmic expansion, by a factor of $a_{\mathrm{final}}\over a_{\mathrm{initial}}.$ Therefore its frequency has been reduced by that same factor, and so we define a redshift factor $z$ in the usual way, $z\,=\,{\omega_{\mathrm{initial}}\over\omega_{\mathrm{final}}}-1\;=\;{a_{\mathrm{final}}\over a_{\mathrm{initial}}}-1.$ If this is so, if the final frequency depends only on the initial and final values of the scale factor, then it seems that no further anisotropy can arise.
-But is this actually the case?
-Is it clear that whatever conditions the photon experiences, in the vast “bulk” of spacetime from decoupling to our location, have no effect on the redshift?
-We need to investigate [^20] The following argument is essentially the solution of the suitably generalised version of Problem 4 in Chapter 5 of [67]..
+But the photons we observe do not emanate from matter during that period : they begin to come from the later time of decoupling (or “recombination”).
+In the interim, the photons cannot stream freely, so it may well be that all signs of the immediate post-Inflationary anisotropy in the redshift field have been lost by the time decoupling occurs.
+So we have to ask : can any torsion which survives to the time immediately after decoupling give rise to still further redshift anisotropy, which might be directly observable?
+By the time of decoupling/recombination, when the CMB photons began their journey towards us, we can assume that the spacetime metric is approximately of the FRW form, with Euclidean spatial sections and a scale factor $a(t)$, where $t$ is the cosmic time defined by using Gaussian normal coordinates constructed using $\nabla^{0},$ the zero-torsion connection defined by $g^{*}$.
+(Then $\xi=\partial_{t}$ when both are evaluated on $\Sigma;$ this relation need not hold off $\Sigma,$ but that will not affect our discussion here.) Let us suppose that the torsion, in particular the bulk torsion, is not yet negligible at this time and for an interval afterwards.
+(We will not assume that it arises from a Weyssenhoff fluid at this time; instead we return to considering torsion of an arbitrary form.)
+We now ask: how does torsion affect the redshift field we actually observe?
+In conventional General Relativistic cosmology, one often sees an argument of the following form.
+When the photon arrives at our location, its wavelength has been “stretched” by the cosmic expansion, by a factor of $a_{\mathrm{final}}\over a_{\mathrm{initial}}.$ Therefore its frequency has been reduced by that same factor, and so if we define a redshift factor $z$ in the usual way, $z\,=\,{\omega_{\mathrm{initial}}\over\omega_{\mathrm{final}}}-1$, then $z$ is given simply by $z\;=\;{a_{\mathrm{final}}\over a_{\mathrm{initial}}}-1.$
+This argument is however based on the assumption that “stretching” is the only effect to which the wavelength is subjected.
+In the presence of “bulk” torsion, that assumption is not correct.
+Separately, the argument is also based on the assumption that the history of the photons does not matter.
+We can explain this second point in a more physical way as follows.
+In a time-dependent spacetime geometry, there is no timelike Killing vector field, and consequently no law of conservation of energy.
+The photon simply loses energy at some points along its path, and gains it at others, depending on the way in which the spatial geometry varies with time; and there is no reason to expect that the gains at one time will compensate for the losses at another.
+Yet the elementary argument we have been discussing insists that, if for example $a_{\mathrm{final}}$ happens to be the same as $a_{\mathrm{initial}}$, then this compensation must happen and be complete and exact.
+One could say that the argument is equivalent to the claim that there is no “memory” of the losses and gains of energy, no “cosmic hysteresis”.
+As we will see shortly, this is actually correct (in General Relativity), but this way of putting the point makes it clear that the standard “elementary” argument is far more subtle than it appears.
+Clearly we need a better understanding of the whole argument [^22] The following discussion is essentially the solution of the suitably generalised version of Problem 4 in Chapter 5 of [81]..
 Using equation (49), we have
 ${d(a\omega)\over d\lambda}\,=\,\omega{da\over d\lambda}\,-\,a\omega^{2}\left(H\,+\,\beta(n,\,n)\,+\,b(n)\right).$ (57)
-Since $H=(da/dt)/a$ in this case, and $\omega=dt/d\lambda,$ we have $a\omega H=da/d\lambda,$ so that there is a cancellation and we obtain
+Since $H=(da/dt)/a$ here, and $\omega=dt/d\lambda,$ we have $a\omega H=da/d\lambda,$ so that there is a cancellation and we obtain
 ${d(a\omega)\over d\lambda}\,=\,-\,a\omega^{2}(\beta(n,\,n)\,+\,b(n)).$ (58)
 It is interesting to note that, in deriving this, we did not actually use the fact that the spatial sections are Euclidean.
-If we set both parts of the bulk torsion equal to zero, we obtain a remarkable result: $a\omega$ has the same value when the photon is received as it had initially.
-This validates the above elementary explanation of the redshift: at the point of observation, the wavelength of the photon has been “stretched” by a factor $a_{\mathrm{final}}/a_{\mathrm{initial}}$, so the frequency (and the energy) has been reduced accordingly, because $a\omega$ is constant along the worldline.
-But what is the physics underlying the constancy of $a\omega?$ Let us try to explain this.
-In a time-dependent spacetime geometry, there is no timelike Killing vector field, and consequently no law of conservation of energy.
-The photon simply loses energy at some points along its path, and gains it at others, depending on the way in which the geometry varies with time; and there is no reason to expect that the gains at one time will compensate for the losses at another.
-If however we imagine for example a zero-torsion cosmology with a periodically fluctuating scale factor, then the fact that $a\omega$ is constant means that the photon always does regain, over a complete cycle, whatever energy it has lost: there is no “memory” of any losses or gains.
-This is the real reason for the fact that the final energy loss depends only on the ratio $a_{\mathrm{final}}/a_{\mathrm{initial}}$ in the usual argument.
-The absence of what might be called “cosmic hysteresis” in zero-torsion cosmology is the real physical meaning of the constancy of $a\omega.$
-When the bulk torsion makes its presence felt, however, the situation is very different. When equation (56) is integrated, we obtain
+If we set both parts of the bulk torsion equal to zero, we obtain a remarkable result : $a\omega$ is constant along the worldline, so it has the same value when the photon is received as it had initially.
+This validates the above elementary explanation of the redshift  in General Relativity: at the point of observation, the wavelength of the photon has indeed been “stretched” by a factor $a_{\mathrm{final}}/a_{\mathrm{initial}}$, so the frequency (and the energy) has been reduced accordingly, without regard to anything that the photon encountered along its worldline; there is no “hysteresis”.
+But this is only true because $a\omega$ is constant along the worldline.
+The tale is very different when “bulk” torsion is present, however. Then, when equation (58) is integrated, we obtain
 ${\omega_{\mathrm{final}}\over\omega_{\mathrm{initial}}}\,=\,{a_{\mathrm{initial}}\over a_{\mathrm{final}}}\,\exp\left(\,-\,\int^{\lambda_{\mathrm{final}}}_{\lambda_{\mathrm{initial}}}\omega\left(\beta(n,\,n)\,+\,b(n)\right)d\lambda\right)\,;$ (59)
-so we find that the ratio of the final frequency to the initial frequency depends on the values taken by the bulk torsion, and on the values of the spatial direction field $n$ (and also on the frequency itself) all along the path —  not just on the initial and final values of the scale factor.
-When the bulk torsion is not zero, the final redshift can no longer be understood simply in terms of the stretching of wavelengths: the loss of energy depends on the geometry of the far-off and long-ago “bulk” of spacetime as the photon passes through it.
+so we find that the ratio of the final frequency to the initial frequency depends on the values taken by the “bulk” torsion all along the path (because $\beta$ and $b$ are evaluated on $n$) —  not just on the initial and final values of the scale factor.
+When the bulk torsion is not zero, the final redshift can no longer be understood simply in terms of the stretching of wavelengths : the loss of energy depends on the geometry of the far-off and long-ago “bulk” of spacetime as the photon passes through it, and this effect could vary for photons passing through different regions of spacetime —  it could be path-dependent.
 In short, bulk torsion gives rise to “cosmic hysteresis”.
-The dependence on $n$, and on the bulk torsion fields if (as is generically the case) they themselves are not isotropic, means that the redshift field in a torsional theory will be anisotropic in general.
-Even if the bulk torsion is negligible currently and was so in the recent past, the hysteresis could still have a potentially detectable anisotropic effect on the final redshift if the bulk torsion was not negligible even for a brief period in the distant past.
-Specifically: if there was an appreciable bulk torsion at and shortly after decoupling, then equation (57) means that it might have had an observable effect on the redshift now, even if the bulk torsion has diluted, to the point of being unobservable, since that time.
-The exact form taken by this anisotropy is difficult to predict, precisely because of the dependence on the values of $\beta,$ $b$, $\omega$, and $n$ along the worldline of the photon; the integral on the right side of equation (56) will be very difficult to evaluate in practice.
-No doubt a numerical investigation would be necessary to determine it.
-But the essential point is clear: in a torsional theory, the redshift field may well be slightly anisotropic.
-We noted, in Section 1, that we proposed to leave to one side other, non-geometric, potential sources of cosmic anisotropy, including the possibility of bulk flows [34, 35].
-We have seen, in our discussion of the Weyssenhoff fluid, that “tilt” arises very naturally in torsional theories, so the two kinds of anisotropy might well occur simultaneously and be intertwined.
-It might be challenging to separate the two effects.
+In torsional inflationary cosmologies, then, the observed redshift field is generically anisotropic, and the extent of the anisotropy depends on time.
+Of course, the manner in which the redshift anisotropy varies with time will depend on the particular theory chosen.
+One can speculate that, in some cases, the “anisotropy field” could fluctuate upwards, perhaps quasi-periodically, before eventually dying out.
+In such a theory one might not be able to argue against claims of observed anisotropy at later times by using the absence of anisotropy at that level in the CMB, and this could open the way to a full theory for observed anisotropies, if they exist.
 9. Conclusion
@@
 We have seen however that, if we introduce torsion into cosmology, then there is a price to be paid : in all but a few special cases, Inflation no longer completely isotropises the early Universe.
-But recent observations show that this may be precisely what we need —  if indeed one or more of the various kinds of cosmic anisotropy should be confirmed observationally, then torsion may be a way to reconcile such observations with Inflation.
-Not only that : if this works, torsion will clearly be by far the most natural way of effecting that reconciliation.
-However, the operative word here is, as it should be, “if ”.
+But recent observations show that this may be precisely what we need: the evidence for anisotropy is hard to ignore [24].
+If one or more of the various kinds of cosmic anisotropy should be confirmed observationally, then torsion may be a way —  or the way —  to reconcile such observations with Inflation.
+However, the operative word here is, as it should be, “may ”.
 We have seen that torsion can, extremely naturally, account for some forms of anisotropy, but by no means all.
-Anisotropy along more than one axis ; anisotropy that involves no rotation ; anisotropy such that the magnitude of the Hubble field is not minimal along the distinguished axis ; anisotropy not accompanied by anomalies in the early value of the Hubble parameter : any of these, if confirmed, might present a serious challenge to our proposed reconciliation of Inflation with anisotropy.
+Anisotropy (at the end of Inflation) along more than one axis, or along one axis but not minimal there ; anisotropy that involves no rotation, or is such that the amount of anisotropy is not correlated with the amount of rotation ; anisotropy not accompanied by anomalies in the early value of the Hubble parameter : any of these, if confirmed, might present a serious challenge to our proposed reconciliation of Inflation with anisotropy.